### v1

# AutoOPE: Automated Off-Policy Estimator Selection

Abstract
The Off-Policy Evaluation (OPE) problem consists of evaluating the performance of counterfactual policies with data collected by another one. This problem is of utmost importance for various application domains, e.g., recommendation systems, medical treatments, and many others. To solve the OPE problem, we resort to estimators, which aim to estimate in the most accurate way possible the performance that the counterfactual policies would have had if they were deployed in place of the logging policy. In the literature, several estimators have been developed, all with different characteristics and theoretical guarantees. Therefore, there is no dominant estimator, and each estimator may be the best one for different OPE problems, depending on the characteristics of the dataset at hand. While the selection of the estimator is a crucial choice for an accurate OPE, this problem has been widely overlooked in the literature. We propose an automated data-driven OPE estimator selection method based on machine learning. In particular, the core idea we propose in this paper is to create several synthetic OPE tasks and use a machine learning model trained to predict the best estimator for those synthetic tasks. We empirically show how our method is able to generalize to unseen tasks and make a better estimator selection compared to a baseline method on several real-world datasets, with a computational cost significantly lower than the one of the baseline.
1 Introduction
In numerous real-world scenarios, decision-making systems are required to interact with dynamic environments, making choices aimed at optimizing specific objectives. These systems are prevalent in domains such as recommender systems [10, 30, 7], information retrieval [22], ad-placement systems [4], or medical treatments [9]. Such decision-making problems can be viewed through the lens of the Contextual Bandit paradigm [43, 19]. For instance, let us consider the recommendation system domain: the recommender (agent) repeatedly observes user profiles (contexts), and aims to recommend items (actions) to users based on the given context to maximize some notion of reward.
A critical challenge arises in evaluating the performance of new policies without deploying them in the real world. This is especially pertinent in scenarios where deploying a suboptimal policy can have significant negative consequences, such as lost revenue in ad placement or user dissatisfaction in recommender systems. Off-Policy Evaluation (OPE) seeks to address this challenge by evaluating the performance of counterfactual policies using data collected from a different, previously deployed policy, often termed the logging policy. The ability to make informed decisions about policy modifications without necessitating real-world deployment is not only cost-effective but also reduces the potential risks associated with untested policies.
However, the effectiveness of OPE largely hinges on the accuracy of the estimator used to predict the performance of the evaluation policies. There is a wide variety of estimators that have been proposed in the literature, each with distinct characteristics and mathematical properties [41]. As an example, a well-known estimator, called Inverse Propensity Scoring (IPS) [29, 34], mitigates the estimation bias by adjusting observed rewards based on the inverse probability assigned by the logging policy to the logged action. In this way, actions that have a low probability under the logging policy receive higher weighting to compensate for their infrequent occurrence in the dataset, and vice versa for actions more likely under the logging policy. Despite IPS’s prominence, numerous other estimators exist, each with unique trade-offs between bias and variance. Given the heterogeneity of real-world datasets and the myriad of scenarios in which OPE is applied, no single estimator is universally optimal. Further details on related work are reported in Appendix A. An estimator may be the best under certain circumstances but perform poorly under others. This variety of estimators, while offering flexibility, also introduces the challenge of Estimator Selection, which is a crucial problem in OPE and has been mostly overlooked in the existing literature. Basically, the Estimator Selection problem consists of the following: we are given an Off-Policy Evaluation task and we have to select the best estimator for the task we have at hand.
In this paper, we address the Estimator Selection problem using an automated data-driven approach. Our main idea consists of exploiting synthetic data. We generate several different synthetic OPE tasks by varying the characteristics of each one. Various types of estimators are applied to these tasks, and their estimates are compared against the ground-truth value of each evaluation policy, which is known due to the synthetic nature of the data. These synthetic OPE tasks are used to train a machine learning model designed to predict estimator’s Mean Squared Errors for the given OPE task. This model can select the estimator with the lowest predicted error, effectively addressing the OPE Estimator Selection challenge. We call our Estimator Selection method Automated Off-Policy Estimator Selection (AutoOPE). The problem of Estimator Selection has been widely overlooked in the literature, and, to the best of our knowledge, only a very recent paper [40] proposed a method for Estimator Selection, named Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF), which is substantially different from our proposal.
In this paper, we propose a novel data-driven approach called Automated Off-Policy Estimator Selection (AutoOPE) to address the Estimator Selection problem in Off-Policy Evaluation. Our key contributions are:
  • We train a machine learning model to predict the performance of estimators on a large number of synthetic OPE tasks with varying characteristics. We share the code for our experiments for researchers and practitioners in the field [^1] We anonymously share the code to reproduce our experiments: https://anonymous.4open.science/r/auto-ope-28F3/.
  • To the best of our knowledge, this is the first work to address the estimator selection problem in OPE using a data-driven approach leveraging only synthetic data, and it is one of the few techniques proposed for Estimator Selection in OPE in general, despite the importance of the problem.
  • We conduct extensive experiments on several datasets that empirically demonstrate the effectiveness of our proposed AutoOPE method in selecting high-performing estimators tailored to the given OPE task, performing better than the baseline with a much reduced computational cost. This represents an important contribution towards automated and adaptive Off-Policy Evaluation.
2 Background
2.1 Contextual Bandit Problem
The Contextual Bandit (CB) problem is an instance of sequential decision-making problems. At a given time instant $t$, an agent observes a context $x_{t}\sim p(\cdot)$ from a space $\mathcal{X}$, and chooses an action $a_{t}$ from a set $\mathcal{A}$, according to a policy $\pi_{b}$, that is a statistical distribution of the actions conditioned on the context ($a_{t}\sim\pi_{b}(\cdot|x_{t})$), called logging policy. After choosing an action $a_{t}$ the agent receives a reward $r_{t}$ from the environment, sampled from $r_{t}\sim p(\cdot|x_{t},a_{t})$. This means that the reward observed is conditioned on both context and action.
A common quantity used to compare the quality of different policies is the policy value, that is defined as the expected value of the reward collected under a certain policy $\pi$: $V(\pi)\coloneqq\mathbb{E}_{p(r_{t}|x_{t},a_{t})\pi(a_{t}|x_{t})p(x_{t})}\left[r_{t}\right]$.
2.2 Off-Policy Evaluation
The Policy Evaluation problem refers to the task of estimating the policy value of a given policy $\pi_{e}$, called evaluation policy, and understanding how it would perform in a given environment. In this paper we focus on Off-Policy Evaluation (OPE). In OPE, we would like to estimate the policy value of the evaluation policy $\pi_{e}$ using data collected under a different logging policy $\pi_{b}$. This logging data is defined as $\mathcal{D}_{b}=\left\{\left(x_{t},a_{t},r_{t},\left\{\pi_{b}(a|x_{t})\right\}_{a\in\mathcal{A}}\right)\right\}_{t}^{n_{b}}$, with $n_{b}$ being the number of tuples in $\mathcal{D}_{b}$, also called number of rounds. We define an OPE task $\mathcal{T}_{\text{ope}}$ as the set composed of the logging dataset and the evaluation policy:
$\mathcal{T}_{\text{ope}}\coloneqq\left\{\mathcal{D}_{b},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}$ (1)
The problem of Off-Policy Evaluation can be tackled with OPE estimators, which can be defined as a statistical function that uses logging data $\mathcal{D}_{b}$ collected under the logging policy $\pi_{b}$ as well as the evaluation policy $\pi_{e}$ computed for each context $x_{t}$, to estimate the ground-truth evaluation policy value $V(\pi_{e})$. We denote an OPE estimator as $\hat{V}(\pi_{e})$.
A common metric used to assess the performance of an estimator is the Mean Squared Error (MSE), defined as $MSE(\hat{V}(\pi_{e}))=\mathbb{E}\left[(\hat{V}(\pi_{e})-V(\pi_{e}))^{2}\right]$.
2.3 Estimator Selection for OPE
The effectiveness of OPE crucially depends on how accurate is the estimator in estimating the policy value of the evaluation policy, usually in terms of how low is its $MSE(\hat{V}(\pi_{e}))$. For this reason, and since many OPE estimators have been proposed in the literature, selecting the best estimator for a given OPE task, i.e., Estimator Selection, is particularly important. Given a finite set $\mathcal{V}$ of different OPE estimators, the Estimator Selection problem consists in finding the best estimator among them.
Since in OPE we do not have any data collected under the evaluation policy, we cannot compute the ground-truth evaluation policy value $V(\pi_{e})$ and therefore the ground-truth MSE of the estimator. A common approach is to approximate it with an estimate, i.e., $\widehat{MSE}(\hat{V}(\pi_{e}))$, and to select the estimator that has the lowest approximate MSE.
2.3.1 Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF)
To the best of our knowledge, only Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF) [40] has tackled the problem of Estimator Selection in the literature. To accurately estimate the MSE of an OPE estimator, the key idea of PAS-IF is to determine a sampling rule $\rho_{\theta}:\mathcal{X}\times\mathcal{A}\rightarrow(0,1)$, that governs a subsampling procedure to divide the logging data $\mathcal{D}_{b}$ into two pseudo datasets: a pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ and a pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$. $\rho_{\theta}(x_{t},a_{t})$ represents the probability of assigning the logging sample $(x_{t},a_{t})$ to $\tilde{\mathcal{D}}_{e}$. In this context, the two pseudo datasets can be viewed as if they were collected by two pseudo policies, $\tilde{\pi}_{b}$ and $\tilde{\pi}_{e}$, induced by $\rho_{\theta}$. The $\rho_{\theta}(x_{t},a_{t})$ is the output of a fully connected neural network, optimized by minimizing the average squared error between the importance ratio $w(x_{t},a_{t})=\pi_{e}(a_{t}|x_{t})/\pi_{b}(a_{t}|x_{t})$ induced by true policies, and $\tilde{w}(x_{t},a_{t})=\tilde{\pi}_{e}(a_{t}|x_{t})/\tilde{\pi}_{b}(a_{t}|x_{t})$ induced by the pseudo policies: this learning phase is called importance fitting. The pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$ is deemed to be collected by the evaluation policy, and in this way it can be used to estimate the evaluation policy value by averaging the rewards in it (we denote this estimate as $\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})$). The pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ is used to compute the candidate OPE estimator prediction ($\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b})$). After that, each estimator MSE can be approximated in this way:
$\widehat{MSE}_{\mbox{PAS-IF}}(\hat{V}(\pi_{e}))\coloneqq(\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})-\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b}))^{2}\ .$
One limitation of PAS-IF is that $\rho_{\theta}(x_{t},a_{t})$ is different for every OPE task, hence requiring to re-train a separate PAS-IF model for each OPE task. A second limitation is that the ability of PAS-IF to build two effective pseudo datasets is constrained by the available training data which may be problematic in scenarios where the data is limited.
3 Automated Off-Policy Estimator Selection
In this section, we present a novel approach for solving the Estimator Selection problem, called Automated Off-Policy Estimator Selection (AutoOPE).
3.1 Theoretical Motivation
AutoOPE is theoretically justified by many findings that show how the performance of an OPE estimator is highly dependent on the characteristics of the OPE task at hand [42, 32]. For instance, [24, 25, 26] theoretically analyze Inverse Propensity Scoring (IPS) [29, 34], one of the most used OPE estimators and the basis for many others. Metelli et al. [24] show that the IPS error is strictly related to the (exponentiated) Rényi divergence between the logging and evaluation policies [^2] The Rényi divergence is a divergence between two probability distributions, and it is defined as $d_{r}(\pi_{e}||\pi_{b})=\log\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\ a\sim\pi_{b}}\left[(\pi_{e}(a|x)/\pi_{b}(a|x))^{2}\right]$. and provide a finite-sample bound on the absolute estimation error of IPS under standard assumptions:
$\left\lvert\hat{V}_{\text{IPS}}(\pi_{e})-V(\pi_{e})\right\rvert\leq\mathcal{O}\left(\sqrt{\frac{\exp(d_{r}(\pi_{e}||\pi_{b}))}{n_{b}}}\right),$
later also proving that this bound is tight [26]. This theoretical finding, along with many others, shows how the performance of an OPE estimator is strictly related to the characteristics of the OPE task, such as the divergence between the evaluation and the logging policy. In other words, we assume that:
$MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})=f^{*}(g(\hat{V},\mathcal{T}_{\text{ope}}))\ ,$
which means that there is a function $f^{*}$ that express the relationship between the real error of an estimator on a given OPE task $MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})$ and the features $g(\hat{V},\mathcal{T}_{\text{ope}})$ of the estimator and the task. For this reason, we propose our zero-shot meta-learning approach [8], which we frame as a supervised learning problem.
Let us assume that we have a large variety of tasks at our disposal $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$, a set of estimators $\mathcal{V}$, and the feature extraction function $g$. Also, let us assume that we have access to the ground-truth MSE. If we have access to these, we could try to find the function that best predicts the MSE for all tasks and estimators, as a standard supervised problem:
$\hat{f}=\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)\ ,$
where $\mathcal{L}$ is a suitable loss. Once we found $\hat{f}$, we can apply it in a zero-shot way whenever we need to perform estimator selection with a new task $\mathcal{T}_{\text{ope}}^{(\text{new})}\notin\mathcal{M}$. Indeed, with $\hat{f}$ we can now estimate the MSE for a given task in this way:
$\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right)\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})).$
Hence, we just need to select the estimator that minimizes the MSE estimated by AutoOPE:
$\hat{V}^{*}=\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right).$
3.2 Practical Implementation
However, this approach is feasible if we have access to several OPE tasks, a feature extraction function, and the ground-truth MSE. For this reason, our approach extensively rely on synthetic data. Specifically:
Dataset
We generate a meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ composed of several different synthetic OPE tasks. In this way, we have access to the ground-truth MSE. Also, since the objective is to obtain a model that is able to generalize in a zero-shot way, we want to train our model with a large number of OPE tasks. This would be unfeasible using real-world OPE data, which is scarce. We indeed generate 250,000 synthetic OPE tasks. Furthermore, to favor generalization of our model, we create OPE tasks with different numerical characteristics. For more details on the synthetic data generation, we refer to Appendix C.
Features
Regarding the features $g$, while theoretical results have proven a relation, for instance, with the exponentiated Rényi divergence, we want to take into account a wide selection of possible features. Hence, we designed a total of 43 features. In particular, we categorize the features into three groups: (i) Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward variance, the number of actions. (ii) Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different statistical distances and divergences. (iii) OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, or uses importance sampling. See Appendix E for their full description.
Estimators
The set $\mathcal{V}$ of candidate estimators include several model-free, model-based and hybrid estimators, as done in [40]. All the estimators used are described in Appendix B.
Model
As a supervised model $\hat{f}$, we use a Random Forest regressor [5] because of its high-quality predictions on tabular data and its speed, both in training and inference. For more details on training and optimization, we refer to Appendix D.
We recap the AutoOPE method in Algorithms 1 (training of the model) and 2 (zero-shot estimator selection).
Algorithm 1 AutoOPE training
0:  Feature generator $g$, estimator set $\mathcal{V}$
0:  MSE estimator $\widehat{MSE}_{\mbox{AutoOPE}}$
1:  Generate meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ (Appendix C)
2:  $\mathcal{L}\leftarrow$ Random Forest regression loss
3:  $\hat{f}\leftarrow\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)$ (Train the random forest)
4:  For any $\hat{V},\mathcal{T}_{\text{ope}}$, define $\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}\right)\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}))$
5:  return  $\widehat{MSE}_{\mbox{AutoOPE}}$
Algorithm 2 AutoOPE estimator selection
0:  New OPE task $\mathcal{T}_{\text{ope}}^{(\text{new})}$, estimator set $\mathcal{V}$, MSE estimator $\widehat{MSE}_{\mbox{AutoOPE}}$
0:  Selected estimator $\hat{V}^{*}$
1:  $\hat{V}^{*}\leftarrow\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right)$
2:  return  $\hat{V}^{*}$
The AutoOPE approach has several advantages. First of all, it is data-driven: AutoOPE tries to learn the shared patterns in several different synthetic OPE tasks with the objective to generalize to new unseen tasks. For this, we use a Random Forest, which is a powerful and fast model for tabular data. Furthermore, it is theory-inspired: it stems from theoretical findings to create a data-driven model that takes in input features such as divergences which are proven to be related to the estimator’s error, but also features that have no theoretical guarantee at the moment. Thanks to this, we can also provide insights that can shed light on future theoretical findings by analyzing the features that had a significant impact on the model’s performance. Finally, it is zero-shot. This means that a practitioner can avoid training for each new estimator selection task, leading to a much faster computational time.
A possible downside of our approach is that we trained AutoOPE on large meta-dataset $\mathcal{M}$ that is synthetic and created a priori, which may not lead to generalization to particular real-world OPE tasks. However, in Section 4, we show that AutoOPE generalizes well to unseen real-world OPE tasks with an extensive experimental evaluation on a large number of real-world OPE tasks.
4 Experiments
In this section, we describe the experimental analysis of our proposed AutoOPE and show it is competitive with the PAS-IF baseline on several real-world tasks, as well as exhibiting a lower computational cost.
4.1 Experimental Settings
PAS-IF Baseline
Our proposed method is compared to the PAS-IF baseline, which has specific hyperparameters. One important drawback of PAS-IF is the high computational cost, see Section 4.3, which makes hyperparameter tuning impractical. [^3] We estimate that applying the same Bayesian Search used for AutoOPE would require at least a month of high-end GPU usage. For this reason we rely on the ad-hoc strategy presented in [40] which tunes the regularization parameter $\lambda$ for each new OPE task via grid search.
AutoOPE
We train the Random Forest model of AutoOPE on the synthetic meta-dataset $\mathcal{M}$ and we optimize its hyperparameters with a Bayesian search. For more details, we refer to Appendix D. In this Section, we only use the trained Random Forest in inference, in a zero-shot way (see Algorithm 2) on new unseen tasks.
Evaluation Metrics for OPE Selection
To evaluate AutoOPE and PAS-IF on the Estimator Selection task we use two metrics: Relative Regret and Spearman’s Rank Correlation Coefficient, as in [40], leveraging the fact that in all experiments we have access to the ground-truth MSE of each estimator. Further details are in Appendix H.
[…]
Spearman’s Rank Correlation Coefficient [33] measures the similarity between two rankings, one based on the MSE predicted by the Estimator Selection method and the other based on the ground-truth MSE.
4.2 Results
Here we present the results we obtained on several public real-world datasets.
4.2.1 Open Bandit Dataset (OBD)
In this real-world experiment, we use the Open Bandit Dataset (OBD) [31], a publicly available dataset originated from a large-scale fashion e-commerce platform. The dataset encompasses three campaigns: ‘ALL’, ‘MEN’, and ‘WOMEN’. In each campaign, the deployed policy is randomly chosen at each round, selecting either the Uniform Random policy or the Bernoulli Thompson Sampling policy.
CB Data
We focus on the ‘ALL’ campaign, and opt to employ the bandit feedback collected under the Uniform Random policy as logging data, while the data collected under the Bernoulli Thompson Sampling policy serves as evaluation data. In order to compute confidence intervals, we perform stratified bootstrap sampling 20 times. Each subsample consists of around 35,000 rounds and constitutes a new OPE task.
Results
Method Rel. Regret ($\downarrow$) Spearman ($\uparrow$)
PAS-IF 4.54 $\pm$ 4.47 0.08 $\pm$ 0.46
[…]
Table 1: Experimental results on Open Bandit Dataset (OBD).
The outcomes of this experiment are shown in Table 1. Our AutoOPE method exhibits a Relative Regret almost 6 times lower compared to PAS-IF, and the Spearman’s Rank Correlation Coefficient is much better than PAS-IF. AutoOPE confirms its stability in its predictions, exhibiting low variance, while PAS-IF has a large variance especially on Relative Regret. This experiment confirms that AutoOPE, while being trained on synthetic data, is able to generalize on data coming from real distributions.
4.2.2 UCI datasets
To further strengthen our empirical evaluation of the generalization capabilities of AutoOPE, in this section we present 40 other real-world experiments conducted on 8 different datasets. In particular, we took 8 datasets from the UCI repository [1], as typically done in the OPE literature [2, 35, 6, 32]. We report an overview of these datasets in Table 2.
$\begin{array}[]{l|llllllll}\hline\cr\text{Dataset}&\text{letter}&\text{optdigits}&\text{page-blocks}&\text{pendigits}&\text{satimage}&\text{vehicle}&\text{yeast}&\text{breast-cancer}\\ \hline\cr\text{Classes}&26&10&5&10&6&4&10&2\\ \text{Sample size}&20000&5620&5473&10992&6435&846&1484&569\\ \hline\cr\end{array}$
Table 2: Characteristics of the considered UCI datasets.
CB Data
To adapt these datasets for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features as context features and the class labels as actions, while the reward is 1 if the current action is the correct class for the given context, 0 vice versa. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using parameters $\alpha_{b}$, $\alpha_{e}$:
$\pi_{b}(a|x)\coloneqq\alpha_{b}\cdot\pi_{det,b}(a|x)+(1-\alpha_{b})\cdot\pi_{u}(a|x)\ \ \ ,$
$\pi_{e}(a|x)\coloneqq\alpha_{e}\cdot\pi_{det,e}(a|x)+(1-\alpha_{e})\cdot\pi_{u}(a|x)\ \ \ ,$
[…]
Results
Figure 1: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Estimator Selection experiments on various UCI datasets (the name of the used dataset is in the title of the corresponding plot). Shaded areas correspond to 95% confidence intervals.
The results are shown in Figure 1. AutoOPE consistently outperforms PAS-IF, demonstrating lower Relative Regret and better Spearman’s Rank Correlation Coefficient in almost all the 40 different experimental configurations. Indeed, AutoOPE outperforms PAS-IF 36 out of 40 configurations in terms of Spearman’s Coefficient, and 37 out of 40 configurations in terms of Relative Regret, even if AutoOPE performs estimator selection in a zero-shot way. Notably, when AutoOPE outperforms PAS-IF, it does it by a large margin. Also, AutoOPE exhibits lower variance in particular for the Spearman’s coefficient, while PAS-IF exhibits very high variance. This comprehensive empirical evaluation and its results are consistent with the OBD experiments and confirm that AutoOPE is able to generalize to unseen OPE tasks with a different distribution compared to the one used for its training, in a zero-shot fashion.
4.3 Computational Time Comparison
Figure 2: Comparison on the computational cost of AutoOPE and PAS-IF for the Estimator Selection task.
Figure 3: Feature importance of the most and least important features for AutoOPE.
One important advantage of our AutoOPE method is that, after the pre-training phase, it is ready-to-use for inference, potentially on many different datasets and evaluation policies. On the other hand, the baseline method PAS-IF must be trained from scratch and perform hyperparameter optimization for every new evaluation policy and dataset. Due to this, the time required to perform the Estimator Selection task is very different. Figure 3 reports a comparison of the time required to perform the Estimator Selection task. As expected, AutoOPE is in the order of 10 times faster compared to PAS-IF, even though PAS-IF performs estimator selection on a GPU, while AutoOPE does it on a CPU. Further details are reported in Appendix G.
4.4 Additional Experiments
Finally, we conducted other additional experiments on more datasets, various ablation studies, as well as an analysis of the most important features of the AutoOPE model.
Feature Importance
A useful benefit of using Random Forest models is that we can compute the feature importance. For this analysis we use the Mean Decrease in Impurity (MDI), meaning that features with higher importance scores are more influential in the model decision-making process. Figure 3 shows a selection of features for the Random Forest model. The full plot is in Figure 4 of Appendix E. To assess the importance of each feature group, let us focus on the 15 most and least important features in the full plot of Figure 4 in Appendix E. The 11 policy-independent features appear to be of limited importance and correspond to 4 of the 15 least important features. Only 1 policy-dependent feature belongs to the first 15 most important features, which is the 4th moment of the logged rewards and is the second most important feature. Policy-dependent features, which are 24, are generally useful for the model, with 10 among the 15 most important features and 6 out of 24 among the 15 least important. The KL divergence is the most important feature, which suggests further theoretical study may be warranted. The 8 OPE estimator features are similarly spread, with 4 among the most important features and 5 among the least important ones. In particular, among the most important OPE estimator features there are features that characterize the type of estimator (if it uses a reward model or importance sampling or both).
The insights gained from this feature importance analysis is an important contribution. In particular, it can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
Additional Experiments and Ablation studies
We conducted additional experiments on estimator selection on synthetic datasets and on the CIFAR dataset [18]. Also, we conducted different ablation studies, investigating the impact of scaling the size of the meta-dataset, the diversity of the meta-dataset, and the features used. We describe all these additional experiments in Appendix F.
5 Conclusions
In this paper, we focused on the Estimator Selection problem in Off-Policy Evaluation. We showed that, despite its significance, it is an overlooked problem in the literature. We proposed AutoOPE, a novel data-driven approach for Estimator Selection, which leverages synthetic OPE tasks in order to learn to generalize the estimator selection capabilities to unseen real-world OPE tasks. Through several experiments, we empirically showed how our method is able to consistently make a better estimator selection when compared to the baseline method on several real-world datasets, outperforming the baseline method by a large margin with a significantly lower computational cost.
For future work, a natural direction to follow would be to try different machine learning models for AutoOPE, in particular for scenarios where one may want further guarantees on the Estimator Selection process. Furthermore, one could try to theoretically analyze the relationship between the features of the AutoOPE meta-dataset and the performance of the various estimators, which was unveiled by the feature importance of the random forest model. Finally, our approach focused on the estimator selection problem in the contextual bandit domain. Our method could be used as a stepping stone towards improving OPE estimator selection in the more general Reinforcement Learning scenario.
Appendix A Related Work
In the realm of Contextual Bandit Off-Policy Evaluation, there exists a rich and diverse landscape of published research aimed at assessing the performance of policies without actually deploying them, avoiding the hazard of selecting action under not valuable or deleterious policies which, depending on the scenario, could be costly and dangerous [10, 21, 30, 17]. Indeed, this area of study has brought contributions to various application domains, including recommender systems [22], advertising [4], and precision medicine [23], reflecting its flexibility in modeling different situations. In this context, AutoOPE emerges as a novel approach, offering a unique solution based on a black-box model that leverages synthetic data and meta-learning to tackle the Estimator Selection problem.
To appreciate the significance of AutoOPE, it is essential to understand the broader context of Contextual Bandit OPE research.
Several pioneering works have laid the foundation for evaluating counterfactual policies in the context of Contextual Bandit. Notable contributions include studies that have introduced fundamental approaches such as the Direct Method (DM) [3], Inverse Propensity Scoring (IPS) [29, 34], and Doubly Robust (DR) [38, 12, 6], which serve as the cornerstone for OPE research, since they define the three main classes of OPE estimators: model-based, model-free and hybrid.
Direct Method (DM) is rooted in machine learning, employing algorithms to estimate policy performance by regressing rewards. Its effectiveness depends on the precision of the reward estimation produced by the learning algorithm. For this reason, its performance may decrease specially in complex environments like industrial recommender systems. DM is susceptible to bias when the model does not produce accurate reward estimations, caused by the inability to correct the distribution shift between the observed data and the counterfactual policy. It is a purely model-based estimator.
Inverse Propensity Scoring (IPS) instead is purely model-free. IPS is unbiased on expectation, thanks to the employment of the notion of importance sampling by statistics, but exhibits high variance, that increases when logging and evaluation policies are significantly different.
Doubly Robust (DR) represents a blend of DM and IPS characteristics, since it is unbiased and generally presents a lower variance compared to IPS. However, its performance still depends on the specific task so it is not necessarily the best option.
A multitude of OPE estimators, has been designed over time as extensions or combinations of DM, IPS and DR [37, 14, 35, 26, 38, 44]. These efforts collectively contribute to advance OPE methodologies.
However, many estimators incorporate hyperparameters to balance the bias-variance trade-off given the Contextual Bandit environment such as [35, 26, 38, 12, 6, 44]. As reported by [32, 31] the hyperparameter tuning strongly affects the prediction quality. To address the issue of hyperparameter optimization [36] and [39] proposed SLOPE and its improved version SLOPE++, based on Lepski’s principle [20], offering improved results compared to naïve tuning procedures. Lepski’s principle allows to define an upper-bound for the Mean Squared Error without delving into the bias estimation, that is the main reason why the task of MSE estimation is challenging. SLOPE dominates all the other methods designed for hyperparameter tuning, being the state of the art technique for this task. Other existing methods like MAGIC [38], or other built-in methods of proposed estimators [35, 44] have tackled hyperparameter selection using MSE estimators, but they offer weaker guarantees compared to SLOPE.
The Estimator Selection problem has emerged as a relatively unexplored but pivotal aspect of OPE research. Empirical studies by [42] and [32] have shown the fundamental role of Estimator Selection on OPE accuracy, highlighting its importance in practical applications.
In this context, the recent PAS-IF (Policy-Adaptive Estimator Selection via Importance Fitting) method [40] emerges as the only baseline for Estimator Selection. PAS-IF uses a neural network to learn a subsampling rule that is used to generate two pseudo-datasets starting from the available logging data. The subsampling rule is learned by optimizing an importance fitting loss that depends on the specific logging and evaluation policies one wishes to evaluate. Once the two pseudo datasets have been generated, PAS-IF allows to estimate the candidate estimators’ Mean Squared Error based on the empirical average of rewards in these pseudo datasets.
Our proposed AutoOPE method, on the other hand, competes with PAS-IF in performance by introducing a novel data-driven approach that leverages meta-learning with synthetic data and a black-box model. AutoOPE offers a versatile approach, allowing the comparison of different estimator classes and the selection of the most accurate one. This innovative approach empirically promises to significantly enhance the accuracy of OPE, reducing an important gap in the Contextual Bandit Off-Policy Evaluation field.
Appendix B Off-policy Estimators Description
This Appendix contains a brief characterization of all the OPE estimators, $\mathcal{V}$, used in this paper. Note that the total number of OPE candidate estimators is 21 which includes 3 model-free estimators and 6 requiring a reward predictors, of which three were included.
B.1 Model-Based Estimators
We include one estimator based exclusively on a reward model.
Direct Method (DM)
The DM estimator [3] employs a supervised machine learning model $\hat{q}$ to estimate the mean reward $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$. The policy value prediction is computed in this way:
$\hat{V}_{DM}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\ \ .$
It is a method that exhibits a small variance in general, but since its performance is strictly dependent on the reward model precision, the bias can be high when the reward estimation is not accurate. It can be a good choice when the complexity of the task is not too high.
B.2 Model-Free Estimators
We include three estimators that do not rely on a reward predictor, but instead are built on the statistical concept of importance sampling.
Inverse Propensity Scoring (IPS)
The IPS estimator [29, 34] employs pure importance sampling to re-weight the average of rewards, accounting for the distribution shift between $\pi_{b}$ and $\pi_{e}$, since the average of rewards are collected under the logging policy, while the goal is to evaluate $\pi_{e}$. The policy value prediction is computed in this way:
$\hat{V}_{IPS}(\pi_{e};\mathcal{T}_{\text{ope}})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}w(x_{t},a_{t})r_{t},\ \ \ \ w(x_{t},a_{t})\coloneqq\frac{\pi_{e}(a_{t}|x_{t})}{\pi_{b}(a_{t}|x_{t})}\ \ .$
IPS is provably unbiased, if the condition of full support holds, and consistent, but it usually exhibits high variance, especially with small sample data or with divergent $\pi_{b},\ \pi_{e}$ (in this case the importance weight $w$ could blow up, leading to instability).
Self-Normalized Inverse Propensity Scoring (SNIPS)
The SNIPS estimator [37, 14] tries to stabilize the IPS estimator normalizing the prediction by the empirical mean of the importance weights. The policy value prediction is computed in this way:
$\hat{V}_{SNIPS}(\pi_{e};\mathcal{T}_{\text{ope}})\coloneqq\frac{\sum_{t}^{n_{b}}w(x_{t},a_{t})r_{t}}{\sum_{t}^{n_{b}}w(x_{t},a_{t})}\ \ .$
It effectively reduces the variance in many situations with respect to IPS, and it is still consistent. However, it is biased.
Subgaussian Importance Sampling Inverse Propensity Scoring (IPS-$\lambda$)
The IPS-$\lambda$ estimator [26] shrink the importance ratio smoothly, with a sub-gaussian parameterized by a value $\lambda$ to reduce the IPS variance. The policy value prediction is computed in this way:
$\hat{V}_{IPS-\lambda}(\pi_{e};\mathcal{T}_{\text{ope}},\lambda,\gamma)\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}w_{\lambda,\gamma}(x_{t},a_{t})r_{t},$
$w_{\lambda,\gamma}(x_{t},a_{t})\coloneqq((1-\lambda)w(x_{t},a_{t})^{\gamma}+\lambda)^{\frac{1}{\gamma}},\ \ \lambda\in\left[0,1\right],\ \gamma\leq 1\ \ .$
The $\lambda,\ \gamma$ hyperparameters balance the bias-variance trade-off of the estimator.
B.3 Hybrid Estimators
We include five hybrid estimators, that leverage both a reward predictor and importance sampling.
Doubly Robust (DR)
The DR estimator [38, 12, 6] employs a reward model $\hat{q}$ like DM, but it improves the reward prediction by $\hat{q}$ using a correction term based on importance sampling. The policy value prediction is computed in this way:
$\hat{V}_{DR}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q})\coloneqq\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(w(x_{t},a_{t})(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right)\ \ .$
It has useful statistical properties. It is unbiased, consistent, and in general has a lower variance with respect to IPS.
Self-Normalized Doubly Robust (SNDR)
The SNDR estimator [38, 37, 14] is a variant of DR that shrinks the weights using the same normalization factor used in SNIPS. The policy value prediction is computed in this way:
$\hat{V}_{SNDR}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q})\coloneqq\sum_{t=1}^{n_{b}}\left(\frac{w(x_{t},a_{t})}{\sum_{j=1}^{n}w(x_{j},a_{j})}(r_{t}-q(x_{t},a_{t}))+\frac{1}{n_{b}}\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right)\ \ .$
As in SNIPS, the normalization factor is used to gain stability, but some bias is introduced.
Subgaussian Importance Sampling Doubly Robust (DR-$\lambda$)
The DR-$\lambda$ estimator [26] is a variant of DR that shrinks the weights using the same smooth subgaussian by IPS-$\lambda$. The policy value prediction is computed in this way:
$\hat{V}_{DR-\lambda}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q},\lambda,\gamma)\coloneqq\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(w_{\lambda,\gamma}(x_{t},a_{t})(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right)\ \ .$
The $\lambda,\ \gamma$ hyperparameters balance the bias-variance trade-off of the estimator.
Doubly Robust with Optimistic Shrinkage (DRos)
The DRos estimator [35] is a variant of DR that reduces variance with an importance weight derived by minimizing the sharp bound of estimator MSE. The policy value prediction is computed in this way:
$\hat{V}_{DRos}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q},\lambda)\coloneqq\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(\frac{\lambda w(x_{t},a_{t})}{w(x_{t},a_{t})^{2}+\lambda}(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right),$
where the $\lambda\geq 0$ hyperparameter balances the bias-variance trade-off of the estimator.
Switch
The Switch estimator [44] is a variant of DR that switches between DM and DR, using DM when the importance ratio is below a threshold specified by the value of $\lambda$, and DR when $w(x_{t},a_{t})$ is above the threshold. The policy value prediction is computed in this way:
$\hat{V}_{\mathrm{switch}}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q},\lambda)\coloneqq$
$\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(w(x_{t},a_{t})\mathbb{I}\left\{w(x_{t},a_{t})\leq\lambda\right\}\cdot(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right),$
where the $\lambda\geq 0$ hyperparameter balances the bias-variance trade-off of the estimator.
Appendix C Meta-Dataset
C.1 Characteristics
As stated in Section 3, AutoOPE method is based on a zero-shot meta-learning approach framed as a supervised learning problem. In particular, the goal of this supervised learning problem is to predict, given an estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$ as input, the $MSE$ of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$.
Hence, to train a Random Forest regressor to predict this MSE, we need a finite set of possible estimators $\mathcal{V}$ such that $\hat{V}_{m}\in\mathcal{V}$ defined a-priori, and a large set of OPE tasks $\mathcal{M}$. The set $\mathcal{M}$ is a collection of synthetically generated OPE tasks, obtained from a data generator function $G$, which we describe in further detail in Section C.2.
The training dataset for the Random Forest is obtained extracting suitable input features from the estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, that the Random Forest can work with. This is done using a feature extraction function $g$. Then, the target is the ground-truth $MSE$ of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$. The training dataset can be expressed as: $\{g(\mathcal{V}\times\mathcal{M}),\ MSE(\mathcal{V}\times\mathcal{M})\}$
The meta-learning dataset $\mathcal{M}$ contains $250,000$ OPE tasks, each associated to all the candidate estimators in $\mathcal{V}$ that are $21$, for a total of $5.2$ million samples. Note that we removed the few trivial OPE tasks which had all rewards equal to 0 or 1.
OPE Estimators
The set $\mathcal{V}$ of candidate estimators includes several model-free, model-based and hybrid estimators, as done in [40]. Estimators hyperparameters, when present, were selected using the SLOPE method [36]. All the estimators used are listed in Table 3, and a brief explanation of each of them is reported in appendix B.
Each of the six estimators with a reward model is employed in conjunction with three reward predictors: Random Forest Classifier and Logistic Regression from the Scikit-Learn [^4] Scikit-Learn available with 3-Clause BSD license: https://github.com/scikit-learn/scikit-learn library [28], and LGBM Classifier from the Python package of LightGBM [^5] LightGBM available with MIT license: https://github.com/microsoft/LightGBM/tree/master [16]. The reward models were fitted using a 3-fold cross-fitting procedure to reduce possible overfitting [27].
Type OPE Estimator
Model-Based Direct Method (DM)
Inverse Propensity Score (IPS)
Model-Free Self-Normalized IPS (SNIPS)
Sub-Gaussian IPS (IPS-$\lambda$)
Doubly Robust (DR)
Self-Normalized DR (SNDR)
Hybrid Sub-Gaussian DR (DR-$\lambda$)
DR with a optimistic shrinkage (DRos)
Switch DR (SwitchDR)
Table 3: List of the OPE estimators used in this paper. See Appendix B for their description.
Features
We compute 43 features for each OPE task that can be categorized into three groups:
  • Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward mean and variance, the number of actions.
  • Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different geometrical and statistical distances.
  • OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, uses importance sampling, employs a reward model, and if so, which one.
The comprehensive list of features, together with their mathematical formulations, is reported in Appendix E.
Target
The target variable is the MSE between the prediction of an estimator $\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(j)})\ \text{with}\ \hat{V}_{m}\in\mathcal{V}$ and the on-policy estimated ground-truth policy value $V(\pi_{e})$. Indeed, together with each synthetic OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, we also generate a ground-truth dataset, denoted as $\mathcal{D}_{gt}^{(j)}\coloneqq\left\{(x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{gt}}$, that is a large CB dataset collected under the evaluation policy $\pi_{e}$.
Realizations
In order to improve the robustness of the data, both the features and MSE are computed as the average of $n_{gen}$ realizations $\mathcal{T}_{ope}^{(j)}(s)$ of each $\mathcal{T}_{ope}^{(j)}$. The features and the target MSE are computed for each candidate estimator and averaged across all realizations to obtain the final values used in the meta-learning dataset.
C.2 Generation
Data Generator
We can formalize the generation of the pair $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})$ as the output of a function $G$ which serves as synthetic data generator, so $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})=G(p_{1}^{(j)},\dots,p_{P}^{(j)})$, with $(p_{1}^{(j)},\dots,p_{P}^{(j)})$ drawn from $\Phi=(\phi_{1}(\cdot),\dots,\phi_{P}(\cdot))$. For each new OPE task, we sample a value for each parameter $p_{i}$ from the respective distribution $\phi_{i}$, and generate the datasets pair. Function $G$ is highly configurable since each parameter has a distribution that can be specified, which allows to represent a wide range of scenarios. Among the parameters $\left\{p_{i}\right\}_{i}^{P}$ that change the OPE task characteristics we include: the evaluation and logging policies, the number of possible actions, the number of rounds in the logging data, the context dimension, the reward distribution.
Below we provide the complete list of these parameters, along with their distributions $\left\{\phi_{i}\right\}_{i}^{P}$, providing insight into the characteristics considered for OPE tasks:
  • $n_{a}$: a positive integer, representing the number of possible actions available in the considered Contextual Bandit scenario. In our generation process, this parameter is distributed as a random uniform in the range $(2,20)$;
    $n_{a}\in\mathbb{N},\ n_{a}\sim\mathcal{U}(\left\{2,3,\dots,20\right\})\ \ .$
  • $n_{b}$: a positive integer number representing the number of observed contexts, i.e., the number of contexts in the data $\mathcal{T}_{\text{ope}}$. In our generation process, this parameter is distributed as a random uniform in the range $(100,8000)$;
    $n_{b}\in\mathbb{N},\ n_{b}\sim\mathcal{U}(\left\{100,101,\dots,8000\right\})\ \ .$
  • $d_{x}$: a positive integer number representing the context vector dimension. In our generation process, this parameter is distributed as a random uniform in the range $(1,10)$;
    $d_{x}\in\mathbb{N},\ d_{x}\sim\mathcal{U}(\left\{1,2,\dots,10\right\})\ \ .$
  • $q$: it can be either a function provided by the generation module of OBP [^6] Open Bandit Pipeline (OBP) is an open-source library specifically developed to establish a standardized pipeline tailored for research in Off-Policy Evaluation. It is released under the Apache 2.0 License (https://github.com/st-tech/zr-obp), or None. This parameter identifies a function that determines how the expected reward is computed given the pair of context and action. In our generation process, this parameter is distributed as a random uniform over the discrete finite space
[…]
    $q\sim\mathcal{U}(\mathbb{F}_{r})\ \ .$
    Given a context vector $x_{t}\in\mathbb{R}^{d_{x}}$ and an action $a_{t}\in\left\{0,1,\dots,n_{a}-1\right\}$, these functions are used to define the mean reward function $q(x_{t},a_{t})=\mathbb{E}\left[r_{t}|x_{t},a_{t}\right]$.
    If $q$ is None, the expected reward is defined as $\mathbb{E}\left[r_{t}|x_{t},a_{t}\right]\sim\mathcal{U}(0,1)$, while if $q$ is not None, the definition is more complex. Firstly we need to introduce the symbol $c_{t}\in\mathbb{R}^{n_{a}}$, that is the $a_{t}$-th canonical basis vector relative to the space $\mathbb{R}^{n_{a}}$. Then we can define
    $q(x_{t},a_{t})\coloneqq\sigma(\tilde{x}_{t}^{T}M_{X,A}\ \tilde{a}_{t}+\theta_{x}^{T}\tilde{x}_{t}+\theta_{a}^{T}\tilde{a}_{t})\ \ ,$ (2)
    with $\sigma(\cdot)$ that represent the sigmoid function, $M_{X,A}\in\mathbb{R}^{p_{x}\times p_{a}},\ \theta_{x}\in\mathbb{R}^{p_{x}}$, and $\theta_{a}\in\mathbb{R}^{p_{a}}$, and:
    • if $q$ is the logistic_reward_function, the symbols $\tilde{x}_{t}\in\mathbb{R}^{p_{x}}$ and $\tilde{a}_{t}\in\mathbb{R}^{p_{a}}$, are two feature matrices consisting of all $p_{x}$ polynomial combinations of the elements of $x_{t}$, and all $p_{a}$ polynomial combinations of the elements of $c_{t}$, with degree less than or equal to $p=1$;
    • if $q$ is the logistic_polynomial_reward_function, the symbols $\tilde{x}_{t}\in\mathbb{R}^{p_{x}}$ and $\tilde{a}_{t}\in\mathbb{R}^{p_{a}}$, are two new feature matrices consisting of all $p_{x}$ polynomial combinations of the elements of $x_{t}$, and all $p_{a}$ polynomial combinations of the elements of $c_{t}$, with degree less than or equal to $p=3$;
    • if $q$ is the logistic_sparse_reward_function, the symbols $\tilde{x}_{t}\in\mathbb{R}^{p_{x}}$ and $\tilde{a}_{t}\in\mathbb{R}^{p_{a}}$, are two new feature matrices consisting of $p_{x}$ polynomial combinations of the elements of $x_{t}$, and $p_{a}$ polynomial combinations of the elements of $c_{t}$, with degree less than or equal to $p=1$, randomly selected among all the polynomial combinations. Here the $p_{x},\ p_{a}$ combinations randomly chosen are the 10% of all the possible polynomial combinations.
  • $\beta_{b}$: a pair of real numbers $\beta_{b}=(\left\{\beta_{b,i}\right\}_{i\in{1,2}})$. Both numbers define one logging policy $\pi_{b,i}$ each. In the Contextual Bandit environment, the first half rounds are drawn under $\pi_{b,1}$, the second half under $\pi_{b,2}$. Each number, together with the outcome of $q$, contributes to the computation of logits values, used to derive the actions probabilities for each logging policy. Each parameter $\beta_{b,i}$, referred to as the inverse temperature parameter, uniquely characterizes a policy, signifying the proximity of policy $\pi_{b,i}$ to the optimal policy. Higher $\beta_{b,i}$ values imply closer proximity, $\beta_{b,i}=0$ means a random policy, and negative values denote a greater divergence from the optimal actions distribution. In this general formulation, in our generation process, this parameter is distributed as a 2-dimensional vector, with each component distributed as a random uniform with range $(-10,10)$.
    $\beta_{b}\in\mathbb{R}^{2},\ \beta_{b}\sim(\mathcal{U}(-10,10),\ \mathcal{U}(-10,10))\ \ .$
    In our data generation, we decided to generate half of the OPE tasks with the parameter $\beta_{b}$ as a unique real number. This means that half of the synthetic OPE tasks have only one logging policy, while the other half is collected by two logging policies.
  • $\beta_{e}$: a real number. Analogously to $\beta_{b}$, it defines the evaluation policy, and contributes to the computation of logits values, used to derive the counterfactual actions probabilities, together with the outcome of $q$. In our generation process, this parameter is distributed as a random uniform with range $(-10,10)$.
    $\beta_{e}\in\mathbb{R},\ \beta_{e}\sim\mathcal{U}(-10,10)\ \ .$
  • $f_{\pi_{b}}$: it can be either a function provided by the generation module of OBP, or None. This parameter governs the logging policy of our synthetic dataset. We define the logging policy in terms of $f_{\pi_{b}}$ as follows: $\pi_{b}(a_{t}|x_{t})\coloneqq\textit{softmax}_{a_{t}}(\beta_{b}\cdot f_{\pi_{b}}(x_{t},a_{t}))$. In our generation process, this parameter is distributed as a random uniform over the discrete finite space $\mathbb{F}_{\pi}$.
[…]
    $f_{\pi_{b}}\sim\mathcal{U}(\mathbb{F}_{\pi})\ \ .$
    If $f_{\pi_{b}}$ is None, then, we set $f_{\pi_{b}}\coloneqq q(x_{t},a_{t})$; while if $f_{\pi_{b}}$ is not None, we use a more complex logging policy. First, we need to introduce the symbol $c_{t}\in\mathbb{R}^{n_{a}}$, that is the $a_{t}$-th canonical basis vector relative to the space $\mathbb{R}^{n_{a}}$. Then, we define
    $f_{\pi_{b}}(x_{t},a_{t})\coloneqq\tilde{x}_{t}^{T}M_{X,A}\ \tilde{a}_{t}+\theta_{a}^{T}\tilde{a}_{t}\ \ ,$ (3)
    with $M_{X,A}\in\mathbb{R}^{p_{x}\times p_{a}},\ \theta_{x}\in\mathbb{R}^{p_{x}}$, and $\theta_{a}\in\mathbb{R}^{p_{a}}$, and:
    • if $f_{\pi_{b}}$ is the linear_behaviour_policy, the symbols $\tilde{x}_{t}\in\mathbb{R}^{p_{x}}$ and $\tilde{a}_{t}\in\mathbb{R}^{p_{a}}$, are two feature matrices consisting of all $p_{x}$ polynomial combinations of the elements of $x_{t}$, and all $p_{a}$ polynomial combinations of the elements of $c_{t}$, with degree less than or equal to $p=1$;
    • if $f_{\pi_{b}}$ is the polynomial_behaviour_policy, the symbols $\tilde{x}_{t}\in\mathbb{R}^{p_{x}}$ and $\tilde{a}_{t}\in\mathbb{R}^{p_{a}}$, are two feature matrices consisting of all $p_{x}$ polynomial combinations of the elements of $x_{t}$, and all $p_{a}$ polynomial combinations of the elements of $c_{t}$, with degree less than or equal to $p=3$.
  • $f_{\pi_{e}}$: like $f_{\pi_{b}}$, it can be either a function provided by the generation module of OBP, or None. This parameter governs the evaluation policy. In our generation process, this parameter is distributed as a random uniform over the discrete finite space $\mathbb{F}_{\pi}$. The functions in $\mathbb{F}_{\pi}$ are described by Equation 3.
[…]
    $f_{\pi_{e}}\sim\mathcal{U}(\mathbb{F}_{\pi})\ \ .$
Apart from these randomly sampled parameters, the synthetic data generator $G$ has additional parameters that are fixed and independent from the current generated OPE task. These include:
  • $n_{data}$: a positive integer determining the number of synthetic OPE task data ($\mathcal{T}_{\text{ope}}$) to generate. In total, we generated $n_{data}=250,000$ different OPE task.
  • $n_{gen}$: a positive integer determining the number of realizations of each $\mathcal{T}_{\text{ope}}$. In our generation process, $n_{gen}=10$.
  • $n_{gt}$: a positive integer regulating the number of rounds in $\mathcal{D}_{gt}$. In our generation process, $n_{gt}=100,000$
OPE Task Generation
Introduced the parameters used in the generation process, the procedure for obtaining the two datasets $(\mathcal{T}_{\text{ope}},\mathcal{D}_{gt})$ can be formally described.
The procedure is outlined by Algorithm 3. Internally to the function Generate, called in the Algorithm 3:
  1. (1)
    $n_{gt}$ vectors $x_{t}$ are independently sampled from a $d_{x}-$dimensional random vector, distributed as a normal with zero mean and identity covariance matrix.
  2. (2)
    Then for each context, we compute the expected reward for all available $n_{a}$ actions. These expected rewards are determined using the synthetic reward function $q$: it is a deterministic function that has the role to generate synthetically $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$.
  3. (3)
    The evaluation policy $\pi_{e}(a_{t}|x_{t})$ is defined by this formula $\pi_{e}(a_{t}|x_{t})=\textit{softmax}_{a_{t}}(\beta_{e}\cdot f_{\pi_{e}}(x_{t},a_{t}))$, and it can be computed for each context $x_{t}$ and each admissible action $a_{t}$.
  4. (4)
    Then actions $a_{t}$ are sampled for each context following policy $\pi_{e}$, and the relative rewards $r_{t}$ for each pair $(x_{t},a_{t})$ are sampled from a Bernoulli distribution with a mean of $q(x_{t},a_{t})$.
Afterwards, within the function GenerateOPE, called in the Algorithm 3:
  1. (1)
    other $n_{b}$ vectors $x_{t}$ are sampled from the same $d_{x}$-dimensional normal vector with zero mean and identity covariance matrix.
  2. (2)
    The logging policy $\pi_{b}$ is computed for each context $x_{t}$ and all the available actions.
  3. (3)
    An action $a_{t}$ is then selected for each context $x_{t}$, following the policy $\pi_{b}$.
  4. (4)
    The correspondent reward $r_{t}$ is produced and collected, sampling from the same Bernoulli distribution we introduced before.
These data are sufficient to formally define $\mathcal{T}_{\text{ope}}$ as in Equation 1.
The procedure to generate $\mathcal{T}_{\text{ope}}$ is repeated $n_{gen}$ times, obtaining $n_{gen}$ realizations of the $\mathcal{T}_{\text{ope}}$ dataset.
Algorithm 3 OPE Data Generator $G$
0:  number of OPE task generations $n_{data}$, number of realizations $n_{gen}$ of each OPE dataset, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$ for the evaluation dataset, space of generation parameters $\Phi$
0:  Collection of OPE datasets and evaluation datasets $\left\{\{\mathcal{T}^{(j)}_{ope}(s)\}_{s}^{n_{gen}},\ \mathcal{D}^{(j)}_{gt}\right\}_{j=1}^{n_{data}}$
1:  for $j=1,\ldots,n_{data}$ do
2:     $(n_{a},n_{b},d_{x},q,\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}})\sim\Phi$
3:     $\mathcal{A}\leftarrow\left\{0,\dots,n_{a}-1\right\}$
4:     $\mathcal{D}^{(j)}_{gt}\leftarrow\mathrm{Generate}(\mathcal{A},d_{x},\beta_{e},f_{\pi_{e}},q,n_{gt})$
5:     for $s=1,\ldots,n_{gen}$ do
6:        $\mathcal{T}^{(j)}_{ope}(s)\leftarrow\mathrm{GenerateOPE}(\mathcal{A},d_{x},\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
7:     end for
8:  end for
Appendix D Supervised Model Training
Supervised Model
As a supervised model, we select the Random Forest regressor provided by the scikit-learn package [^7] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html.
Data Split
In order to train and evaluate the AutoOPE supervised model, we split the meta-learning dataset in 80% for optimization and training, and 20% for testing, so that the testing data never contains OPE tasks the supervised model was trained on.
Hyperparameter Optimization
AutoOPE uses a Random Forest as a supervised model. In order to select the best hyperparameters we rely on a Bayesian Search optimizing the 5-fold cross-validation Regret between the predicted MSE of the OPE estimator selected by AutoOPE, $\hat{V}_{\hat{m}}$, and the best estimator according to the MSE on the ground-truth dataset. We define the Regret as:
$\text{Regret}\coloneqq MSE(\hat{V}_{\hat{m}}(\pi_{e}))-MSE(\hat{V}_{m^{*}}(\pi_{e}))\ ,$ (4)
This means that at each step of the Bayesian Search, the training dataset is split into five folds, and a Random Forest regression model (with fixed hyperparameters) is fitted for each of the possible five subsets of four data folds, resulting in five models in total. Each model is then evaluated on the remaining out-of-bag fold in terms of Regret.
The Regret values are averaged over the different five models, and the 5-fold cross-validation metric is obtained, that is the function optimized by the Bayesian Search.
The Bayesian Search is implemented on top of the Scikit-Optimize [^8] Scikit-Optimize available with 3-Clause BSD license: https://github.com/scikit-optimize/scikit-optimize/tree/master library, and involves Bayesian optimization using Gaussian Processes to approximate the objective function, making the assumption that function points are distributed as a multivariate Gaussian. The Bayesian Optimization is executed on the training data, with 100 calls made to optimize the hyperparameters.
Type Hyperparameter Range Prior
Integer n_estimator [200] -
max_depth (1, 100) uniform
min_samples_split (2, 50) uniform
min_samples_leaf (1, 50) uniform
Real max_samples (0.01, 1.00) uniform
max_features (0.10, 1.00) uniform
Categorical criterion ["squared_error"] -
oob_score [True] -
Table 4: Hyperparameters Search Space for AutoOPE Random Forest model Hyperparameters Tuning via Bayesian Search
Appendix E Features Design
This Appendix mathematically describes the input features for the Random Forest in AutoOPE. They are:
  • Policy-Independent features:
    • Number of logging rounds, i.e., $n_{b}$.
    • Number of different possible actions, i.e., $n_{a}$.
    • Number of deficient actions in logging data, i.e., $n_{def}\coloneqq\mid\left\{a\in\mathcal{A}:a\notin\left\{a_{t}\right\}_{t}^{n_{b}}\right\}\mid$.
    • Context vector dimension, i.e., $d_{x}$.
    • Variance of logging selected actions, i.e., $\sigma^{2}(a_{t})\coloneqq Var[a_{t}]$.
    • Mean of logging rewards, i.e., $\bar{r}_{t}\coloneqq\mathbb{E}[r_{t}]$.
    • Standard deviation of logging rewards, i.e., $\sigma(r_{t})=\sqrt{Var[r_{t}]}$.
    • Skewness of logging rewards, i.e.,
      $G_{1}(r_{t})\coloneqq\mathbb{E}\left[\left(\frac{r_{t}-\mathbb{E}[r_{t}]}{\sqrt{Var[r_{t}]}}\right)^{3}\right].$
    • Kurtosis of logging rewards, i.e.,
      $G_{2}(r_{t})\coloneqq\mathbb{E}\left[\left(\frac{r_{t}-\mathbb{E}[r_{t}]}{\sqrt{Var[r_{t}]}}\right)^{4}\right].$
    • Average variance of context components, i.e.,
      $\bar{\sigma}^{2}(x_{t})\coloneqq\sum_{j}^{d_{x}}Var[x_{t,j}]\ \ .$
  • Policy-Dependent features:
    • Max of $\pi_{b}$ average, over contexts, i.e., $\text{max}\ \bar{\pi}_{b}\coloneqq\text{max}_{i}\ \frac{1}{n_{b}}\sum_{t}^{n_{b}}\pi_{b}(a_{i}|x_{t})$.
    • Min of $\pi_{b}$ average, over contexts, i.e., $\text{min}\ \bar{\pi}_{b}\coloneqq\text{min}_{i}\ \frac{1}{n_{b}}\sum_{t}^{n_{b}}\pi_{b}(a_{i}|x_{t})$.
    • Max of $\pi_{e}$ average, over contexts, i.e., $\text{max}\ \bar{\pi}_{e}\coloneqq\text{max}_{i}\ \frac{1}{n_{b}}\sum_{t}^{n_{b}}\pi_{e}(a_{i}|x_{t})$.
    • Min of $\pi_{e}$ average, over contexts, i.e., $\text{min}\ \bar{\pi}_{e}\coloneqq\text{min}_{i}\ \frac{1}{n_{b}}\sum_{t}^{n_{b}}\pi_{e}(a_{i}|x_{t})$.
    • Max importance weight, i.e.,
      $\text{max}\ w(\pi_{b},\pi_{e})\coloneqq\text{max}_{t}\frac{\pi_{b}(a_{t}|x_{t})}{\pi_{e}(a_{t}|x_{t})}\ \ .$
    • Average inverse importance weight, i.e.,
      $\bar{w}(\pi_{b},\pi_{e})\coloneqq\mathbb{E}\left[\frac{\pi_{e}(a_{t}|x_{t})}{\pi_{b}(a_{t}|x_{t})}\right]\ \ .$
    • Number of importance weight clipped with a threshold $\lambda=10$, i.e.,
      #$\#w_{10}(\pi_{b},\pi_{e})\coloneqq\left\lvert\left\{\forall(x_{t},a_{t})\in\mathcal{T}_{\text{ope}},\ w(\pi_{b},\pi_{e})=\frac{\pi_{e}(a_{t}|x_{t})}{\pi_{b}(a_{t}|x_{t})}:w(\pi_{b},\pi_{e})>10\right\}\right\rvert\ \ .$
    • Total variation distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{TV}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\frac{1}{2}\cdot\sum_{i}^{n_{a}}\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{t}|x_{t})\right\rvert\ \ .$
    • Neyman chi-squared distance between $\pi_{b},\pi_{e}$, which we included because it is equivalent (up to a constant) to the exponentiated Rényi divergence (shown to be an important data characteristic by Metelli et al. [24]):
      $\text{Neyman}(\pi_{b}||\pi_{e})=\exp(d_{r}(\pi_{e}||\pi_{b}))-1\ \ .$
      We compute it as:
      $\text{Neyman}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\frac{(\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t}))^{2}}{\pi_{b}(a_{i}|x_{t})}\ \ .$
    • Pearson chi-squared distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{P}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\frac{(\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t}))^{2}}{\pi_{e}(a_{i}|x_{t})}\ \ .$
    • Inner product distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Inner}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\pi_{b}(a_{i}|x_{t})\cdot\pi_{e}(a_{i}|x_{t})\ \ .$
    • Chebychev distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Chebyshev}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\text{max}_{i}\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t})\right\rvert\ \ .$
    • Divergence between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Divergence}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}2\cdot\sum_{i}^{n_{a}}\frac{(\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t}))^{2}}{(\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t}))^{2}}\ \ .$
    • Canberra distance metric between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Canberra}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\frac{\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t})\right\rvert}{\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t})}\ \ .$
    • K-divergence between $\pi_{b},\pi_{e}$, i.e.,
      $\text{K}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\pi_{b}(a_{i}|x_{t})\cdot\log\left(\frac{2\cdot\pi_{b}(a_{i}|x_{t})}{\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t})}\right)\ \ .$
    • K-divergence between $\pi_{e},\pi_{b}$, note that the policies are swapped, i.e.,
      $\text{K}(\pi_{e}||\pi_{b})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\pi_{e}(a_{i}|x_{t})\cdot\log\left(\frac{2\cdot\pi_{e}(a_{i}|x_{t})}{\pi_{e}(a_{i}|x_{t})+\pi_{b}(a_{i}|x_{t})}\right)\ \ .$
    • Jensen-Shannon distance between $\pi_{b},\pi_{e}$, i.e.,
      $\begin{split}\text{JS}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\frac{1}{2}\cdot\left[\sum_{i}^{n_{a}}\pi_{b}(a_{i}|x_{t})\cdot\log\left(\frac{2\cdot\pi_{b}(a_{i}|x_{t})}{\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t})}\right)\right.\\ +\left.\sum_{i}^{n_{a}}\pi_{e}(a_{i}|x_{t})\cdot\log\left(\frac{2\cdot\pi_{e}(a_{i}|x_{t})}{\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t})}\right)\right]\ \ .\end{split}$
    • Kullback-Leibler divergence between $\pi_{b},\pi_{e}$, i.e.,
      $\text{KL}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\pi_{b}(a_{i}|x_{t})\cdot\log\left(\frac{\pi_{b}(a_{i}|x_{t})}{\pi_{e}(a_{i}|x_{t})}\right)\ \ .$
    • Kullback-Leibler divergence between $\pi_{e},\pi_{b}$, note that the policies are swapped, i.e.,
      $\text{KL}(\pi_{e}||\pi_{b})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\pi_{e}(a_{i}|x_{t})\cdot\log\left(\frac{\pi_{e}(a_{i}|x_{t})}{\pi_{b}(a_{i}|x_{t})}\right)\ \ .$
    • Kumar-Johnson distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{KJ}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\frac{\left((\pi_{b}(a_{i}|x_{t})^{2}-\pi_{e}(a_{i}|x_{t})^{2}\right)^{2}}{2\cdot\sqrt{(\pi_{b}(a_{i}|x_{t})\cdot\pi_{e}(a_{i}|x_{t}))^{3}}}\ \ .$
    • Additive symmetric chi-squared distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Additive}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\frac{(\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t}))^{2}\cdot(\pi_{b}(a_{i}|x_{t})+\pi_{e}(a_{i}|x_{t}))}{\pi_{b}(a_{i}|x_{t})\cdot\pi_{e}(a_{i}|x_{t})}\ \ .$
    • Euclidian distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Euclidian}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sqrt{\sum_{i}^{n_{a}}\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t})\right\rvert^{2}}\ \ .$
    • Kulczynski similarity between $\pi_{b},\pi_{e}$, i.e.,
      $\text{Kulczynski}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\frac{\sum_{i}^{n_{a}}\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t})\right\rvert}{\sum_{i}^{n_{a}}\text{min}(\pi_{b}(a_{i}|x_{t}),\ \pi_{e}(a_{i}|x_{t}))}\ \ .$
    • City-block distance between $\pi_{b},\pi_{e}$, i.e.,
      $\text{CB}(\pi_{b}||\pi_{e})\coloneqq\frac{1}{n_{b}}\sum_{t}^{n_{b}}\sum_{i}^{n_{a}}\left\lvert\pi_{b}(a_{i}|x_{t})-\pi_{e}(a_{i}|x_{t})\right\rvert\ \ .$
  • OPE estimators features:
    • Whether the estimator is a Self-Normalized estimator or not,
    • Whether the estimator uses Importance Sampling or not,
    • Whether the estimator uses a Machine Learning reward model or not,
    • Whether the estimator is a Sub-Gaussian estimator or not,
    • Whether the estimator is a Shrinked estimator or not,
    • Whether the estimator is a Switch estimator or not,
    • Whether the estimator uses a Random Forest Classifier as reward model or not,
    • Whether the estimator uses a LGBM Classifier as reward model or not,
    • Whether the estimator uses a Logistic Regression as reward model or not.
In Figure 4 is reported the complete Feature Importance plot, measured with Mean Decrease in Impurity (MDI).
Figure 4: Plot of the Feature Importance measured with Mean Decrease in Impurity (MDI). The most important features are at the top.
Appendix F Additional Experiments and Ablation Studies
F.1 Scaling Experiment
We performed a scaling experiment of AutoOPE by varying the size of the meta-learning dataset $\mathcal{M}$. We take 20% of the OPE tasks in $\mathcal{M}$ as the test set. To investigate the scaling behavior, we train and optimize AutoOPE on different sub-samples of the remaining 80% of the OPE tasks in $\mathcal{M}$. The results are reported in Figure 5 and show that the loss (i.e., the Regret) of AutoOPE follows a scaling behavior, which is particularly popular in the recent Deep Learning literature [15, 11]. By increasing the training data, we observe a decrease in the Regret approximately following a power law in the form: $L=A+\frac{B}{D^{\alpha}}$, where $L$ is the loss measured with the Regret, $D$ is the meta-learning dataset size, and $A$, $B$, and $\alpha$ are scalar coefficients. While we selected the largest meta-learning dataset for the experimental analysis of this paper, the results in Figure 5 show that the dataset size can be adjusted based on the trade-off between computational resources and model performance for the scenario of interest.
Figure 5: Scaling experiment by varying the meta-learning dataset size.
F.2 Additional Ablation Studies
Features
To understand the impact of including different feature numbers and types on AutoOPE’s performance, we conducted the following ablation study. We group the features in the 3 sets defined in the manuscript:
  • policies-dependent;
  • policies-independent;
  • estimators-dependent.
We train 3 models (one for each group) in the same way outlined in Appendix D, and we test them on the same unseen test set used for the scaling experiment. We show the results in Table 5.
$\begin{array}[]{l|l|l}\hline\cr\text{Features subset}&\text{Rel. Regret}(\downarrow)&\text{Spearman}(\uparrow)\\ \hline\cr\text{Estimators-dependent}&13.58&0.23\\ \text{Policies-independent}&5.22&0.35\\ \text{Policies-dependent}&5.14&0.43\\ \text{All}&\mathbf{2.41}&\mathbf{0.48}\\ \hline\cr\end{array}$
Table 5: Ablation study on the features of the model.
From these results, we can see the impact of each feature group: using only estimator-dependent features gives poor results, policies-dependent features are the best individual group of features, and using all the features combined is beneficial. This applies to both regret and Spearman.
Dataset diversity
In order to understand the impact of the meta-dataset characteristics on the performance of AutoOPE, we performed the following ablation study. Here, we simulate different generation methods by subsampling the complete dataset $\mathcal{M}$. For this study, we perform two experiments:
  • First, we keep only the OPE tasks with 5 actions or less and train only on these. This experiment should highlight the importance of having tasks with a higher number of actions in the meta-dataset. To compare this generation method with the one used in the paper, we train a model on a random subset of $\mathcal{M}$ with the same size. The results obtained are reported in Table 6.
  • In the second experiment, we only keep the tasks with a KL-divergence between the policies $\leq 0.1$. This experiment should highlight the importance of having tasks where the policies are different in the meta-dataset. To compare this generation method with the one used in the paper, we train a model on a random subset of $\mathcal{M}$ with the same size. The results obtained are reported in Table 7.
$\begin{array}[]{l|l|l}\hline\cr\text{Generation Methods}&\text{Rel. Regret}(\downarrow)&\text{Spearman}(\uparrow)\\ \hline\cr|A|\leq 5&7.39&0.33\\ \text{All Tasks}&\mathbf{3.66}&\mathbf{0.45}\\ \hline\cr\end{array}$
Table 6: Ablation study on the number of actions in the dataset.
$\begin{array}[]{l|l|l}\hline\cr\text{Generation Methods}&\text{Rel. Regret}(\downarrow)&\text{Spearman}(\uparrow)\\ \hline\cr\text{KL}\leq 0.1&117.48&0.15\\ \text{All Tasks}&\mathbf{16.73}&\mathbf{0.43}\\ \hline\cr\end{array}$
Table 7: Ablation study on the type of policies in the dataset.
From these findings, we can see how a more diverse generation method is beneficial and improves performance compared to a generation method that generates tasks with a low number of actions or with similar logging and evaluation policies.
F.3 Additional Experiments on CIFAR-10
The CIFAR-10 dataset [18] is a widely-used benchmark for image classification tasks, with mutually exclusive labeling.
CB Data
To adapt the CIFAR-10 dataset for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features (i.e., images pixels scaled in $(0,1)$) as context features and the class labels as actions. Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies. In order to generate the policies, the second data split is used to train two Logistic Regression classifiers. Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0. To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using using parameters $\alpha_{b}$, $\alpha_{e}$:
$\pi_{b}(a|x)\coloneqq\alpha_{b}\cdot\pi_{det,b}(a|x)+(1-\alpha_{b})\cdot\pi_{u}(a|x)\ \ \ ,$
$\pi_{e}(a|x)\coloneqq\alpha_{e}\cdot\pi_{det,e}(a|x)+(1-\alpha_{e})\cdot\pi_{u}(a|x)\ \ \ ,$
[…]
Results
Figure 6: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection of experiments on CIFAR-10 dataset. Shaded areas correspond to 95% confidence intervals.
In order to be able to compute the variance, we perform 20 bootstrap samples of the logging data and we perform the experimental tests on each of them.
The results are shown in Figure 6. AutoOPE consistently outperforms PAS-IF, demonstrating lower Relative Regret and better Spearman’s Rank Correlation Coefficient. Notably, AutoOPE exhibits a much lower variance in particular for the Relative Regret, while PAS-IF exhibits very high variance. These results are consistent with the experiments showed in Section 4 and confirm that AutoOPE is able to generalize to unseen OPE tasks with a different distribution compared to the one used for its training, in a zero-shot scenario.
F.4 Additional Synthetic Experiments
Figure 7: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 1. Shaded areas correspond to 95% confidence intervals.
Figure 8: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 2. Shaded areas correspond to 95% confidence intervals.
In this section we describe all the experiments on synthetic data, following the experimental pipeline of [40].
CB Data
The synthetic experiments are based on two different logging policies, Logging 1 and Logging 2. For each of them 21 OPE tasks with different evaluation policies are defined via a parameter $\beta_{e}$.
The synthetic logging data is generated using the following procedure and distributions (same procedure used in [40]):
  • Context vectors are generated as 10-dimensional vectors, sampled from a multivariate normal distribution with zero mean and identity covariance matrix.
  • A total of 10 possible actions is considered.
  • For each context, we compute the expected reward $\mathbb{E}\left[r(x_{t},a_{t})\right]$, for all available actions. These expected rewards are determined using a synthetic reward function from the ‘dataset’ module of OBP, specifically the ‘logistic_reward_function’, described by Equation 2.
The logging policies $\pi_{b}(a_{t}|x_{t})$ in these experiments are defined as a combination of two policies, each represented as $\pi_{i}(a_{t}|x_{t})=\textit{softmax}_{a_{t}}(\beta_{i}\cdot q(x_{t},a_{t}))$, with $i\in\left\{1,2\right\}$: the first half of rounds is generated under $\pi_{1}$, and the second half under $\pi_{2}$. In total, $n_{b,i}=1,000$ contexts are drawn for each policy $\pi_{i}$, resulting in a total number of logging rounds $n_{b}=2000$.
We compute the logging distribution $\pi_{b}(a_{t}|x_{t})$ of all actions, for each context $x_{t}$ sampled, and afterwards, we randomly sample an action $a_{t}$ for each $x_{t}$, according to this distribution. The actual rewards for these pairs $(x_{t},a_{t})$ are sampled from a Bernoulli distribution with a mean of $q(x_{t},a_{t})$. This process yields a series of context-action-reward triplets $\left\{(x_{t},a_{t},r_{t})\right\}_{t}^{n_{b}}$. These triplets, together with the action distribution $\pi_{b}$ computed for all the $x_{t}$, constitute the logging dataset $\mathcal{D}_{b}=\left\{(x_{t},a_{t},r_{t},\left\{\pi_{b}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{b}}$.
The evaluation policy $\pi_{e}$ is described using the same expression as the logging policies, namely $\pi_{e}(a_{t}|x_{t})=\text{softmax}_{a_{t}}(\beta_{e}\cdot q(x_{t},a_{t}))$, and the counterfactual action distribution is computed for each context $x_{t}$ in $\mathcal{D}_{b}$. It can be noticed that, with this evaluation policy definition, the $\pi_{e}$ is uniquely determined by the parameter $\beta_{e}$. At the end, the complete OPE dataset generated for a synthetic experiment is described as $\mathcal{T}_{\text{ope}}\coloneqq\left\{\mathcal{D}_{b},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$.
As already said, we divide the synthetic experiments into two main settings: one named Logging 1, where we set $\beta_{1}=2$ and $\beta_{2}=-2$, and the other called Logging 2, where $\beta_{1}=3$ and $\beta_{2}=7$, in line with the experimental analysis performed by Udagawa et al. [40].
By varying $\beta_{e}$ values in a range from -10 to 10 for both Logging 1 and 2, we consider 21 different OPE experiments in each setting.
To provide precise performance assessments, in all the 21 experiments of both Logging 1 and 2 we generate $n_{data}=100$ realizations of the logging dataset, as done in the procedure outlined in [40]. The performance metrics are subsequently averaged over all realizations to yield more robust performance estimates with their confidence intervals.
Results
The results obtained from Logging 1 are reported in Figure 8 while those of Logging 2 in Figure 8. Since these synthetic experiments are the same of [40], we use their optimal values for the other hyperparameters such as the network structure, regularization, learning rate. In both cases our proposed AutoOPE outperforms PAS-IF across almost all evaluation policies in terms of Spearman’s rank correlation coefficient. In terms of Relative Regret, AutoOPE performs similar to PAS-IF in the first half settings of the Logging 1 experiments. However, for high values of $\beta_{e}$, PAS-IF slightly outperform AutoOPE. In the Logging 2 set of experiments, we see that for all the $\beta_{e}\geq 0$ AutoOPE outperforms PAS-IF, but PAS-IF performs better in the settings with $\beta_{e}<0$. The fact that AutoOPE exhibits better Spearman’s Rank Correlation indicates that, while both methods predict similar values for the best estimator, AutoOPE is more consistent in ranking all of them.
Appendix G Infrastructure and Computational Time Comparison
For the experiments conducted in this paper, we employed different instances from AWS EC2, to match different needs of resources (GPUs, CPUs) according to the different experiments. The Operating System Ubuntu 20.04 was selected for all the tests.
We used the c6i.32xlarge instance with 128 cores and 256 GB RAM for the synthetic dataset generation explained in Appendix C.2. This instance allows to terminate the whole generation process in about 35 hours.
For the AutoOPE model optimization, training and evaluation instead, additional RAM is needed, since some optimization iterations dive into the training of many and very deep decision trees. The instance used in this case is the m6i.32xlarge instance, with 128 cores and 512 GB RAM. The whole process last approximately 20 hours.
Instead, the experiments described in Section 4 require some hardware accelerator to speed-up the training of the neural network used by PAS-IF. Since the training is very demanding, we adopted the powerful g6.12xlarge instance, that owns four GPUs with 24GB of dedicated memory each, 48 CPUs and 192GB of RAM. Additionally, the PAS-IF code has undergone modifications to enable parallel computation wherever feasible, thereby reducing computational overhead during experimentation. In these experiments, it is apparent the high computational cost of PAS-IF, as shown in Figure 9.
Figure 9: Comparison on the computational cost of AutoOPE and PAS-IF for the Estimator Selection task.
We conduct an additional experiment to test the two Estimator Selection methods on one single synthetic OPE task, so that the maximum level of parallelization could be reached during the execution of PAS-IF. Another factor that affects the time performance is the logging data size, so we perform this experiment different times, varying the logging dataset size. Concerning AutoOPE instead, the model loading time is not negligible, so we test the method under two conditions: in the first condition the black-box model is already present in cache, while in the second the model need to be loaded from disk. The results are reported in Figure 10 and show that even considering the parallelized PAS-IF version and the model loading time for AutoOPE, the time required by PAS-IF to perform the Estimator Selection task exceeds the time needed for AutoOPE to load the model from disk and to perform the same Estimator Selection task. This difference is even higher for larger data sizes.
Figure 10: Computational effort of PAS-IF and AutoOPE in one single synthetic OPE task on g6.12xlarge AWS instance, with varying logging dataset dimension. The AutoOPE is tested on two conditions, with the model loaded from cache, or from disk.
Appendix H Ground-Truth Evaluation Policy Value Estimation
H.1 Open Bandit Dataset
With Open Bandit Dataset, the estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$, can be computed using as ground-truth policy value estimate of $\pi_{e}$, the average of the counterfactual rewards of the evaluation dataset $\mathcal{D}_{gt}$ collected under the evaluation policy $\pi_{e}$. In the experiment outlined in Section 4.2.1, $\pi_{e}$ is the Bernoulli Thompson Sampling policy. To ensure accurate ground-truth policy-value estimation, $\mathcal{D}_{gt}\coloneqq\left\{x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{gt}}$ needs to contain a substantial number of rounds, as in this case, in which $n_{gt}=12,357,200$. The following expression describes the estimated ground-truth MSE of an estimator:
$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{gt}}\cdot\sum_{r_{t}\in\mathcal{D}_{gt}}r_{t}\right)^{2}$ (5)
H.2 UCI Datasets and CIFAR-10
In the experiment reported in Appendix F.3 and in the experiments in Section 4.2.2, we were able to compute the policy value of $\pi_{e}$ in the same way, since we use classification datasets. The policy value $\pi_{e}$ is computed as a true counterfactual expected reward on the logging dataset: in fact, the reward associated with each action and context in the logging data is deterministic (the reward is 1 for the correct class label/action relative to the context, and 0 for all other labels/actions). In this experiment the evaluation dataset can be defined as $\mathcal{D}_{gt}\coloneqq\left\{x_{t},\left\{r_{t,v}\right\}_{v}^{|\mathcal{A}|},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$. The estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with CIFAR-10 and UCI data (transformed in Contextual Bandit data as described in Appendix F.3 and in Section 4.2.2), is computed as described by the following expression:
$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{b}}\cdot\sum_{t}^{n_{b}}\sum_{v}^{|\mathcal{A}|}\pi_{e}(a_{v}|x_{t})\cdot r_{t,v}\right)^{2}$ (6)
H.3 Synthetic Data Experiments
The procedure to estimate the ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with synthetic data, leverages on the possibility to produce synthetically a really big number of rounds under $\pi_{e}$. The evaluation data under policy $\pi_{e}$ are generated obtaining a dataset of this type: $\mathcal{D}_{gt}\coloneqq\left\{x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{gt}}$, and this process is repeated different times. Furthermore, owing to the availability of expected rewards for each context-action pair in synthetic data, the policy value is computed as the weighted average of expected rewards, weighted by the action distribution. The whole procedure is described by Algorithm 4.
Algorithm 4 Ground-Truth MSE Estimation of OPE Estimators in Synthetic Data
0:  number of data generations $n_{e}$, $\beta_{e}$ and a synthetic policy evaluation function $f_{\pi_{e}}$ to define $\pi_{e}$, $\beta_{b}$ and a synthetic policy evaluation function $f_{\pi_{b}}$ to define $\pi_{b}$, $n_{b}$ rounds to sample under $\pi_{b}$, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$, synthetic reward function $q$, OPE estimator $\hat{V}_{m}$, space of available action $\mathcal{A}$
0:  Estimated ground-truth policy value of $\pi_{e},\ \hat{V}_{gt}(\pi_{e})$
1:  for $s=1,\ldots,n_{e}$ do
2:     $\mathcal{D}^{(s)}_{gt}\coloneqq\left\{x_{t},a_{t},r_{t},\left\{f_{\pi_{e}}(x_{t},a;s),q(x_{t},a;s)\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{gt}}\leftarrow\mathrm{Generate}(\beta_{e},f_{\pi_{e}},q,n_{gt},s)$
3:      $\mathcal{T}^{(s)}_{ope}\coloneqq\left\{x_{t},a_{t},r_{t},\left\{f_{\pi_{b}}(x_{t},a;s),f_{\pi_{e}}(x_{t},a;s),q(x_{t},a;s)\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}\leftarrow$ $\mathrm{Generate}(\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
4:  end for
5:  $MSE(\hat{V}_{m}(\pi_{e}))\leftarrow\frac{1}{n_{e}}\cdot\sum_{s}^{n_{e}}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(s)})-\frac{1}{n_{gt}}\cdot\sum_{t}^{n_{gt}}\sum_{a_{j}\in\mathcal{A}}f_{\pi_{e}}(x_{t},a_{j};s)\cdot q(x_{t},a_{j};s))^{2}$
In Appendix F.4 are specified mostly of the input parameters for Logging 1 and 2 used in Algorithm 4, a part from $n_{e},\ n_{gt}$, set with values $n_{e}=100,\ n_{gt}=1,000,000$.
Appendix I Broader impact statement
With an improved performance in Estimator Selection, we improve the quality of OPE. This may be beneficial to society, allowing decision-makers to test policies that could be potentially dangerous without deploying them on a real system with improved reliability. Furthermore, using a pre-trained and not computationally expensive model can be beneficial in terms of environmental impacts.

### v2

@@
-AutoOPE: Automated Off-Policy Estimator Selection
+Automated Off-Policy Estimator Selection via Supervised Learning
 Abstract
 The Off-Policy Evaluation (OPE) problem consists of evaluating the performance of counterfactual policies with data collected by another one.
-This problem is of utmost importance for various application domains, e.g., recommendation systems, medical treatments, and many others.
 To solve the OPE problem, we resort to estimators, which aim to estimate in the most accurate way possible the performance that the counterfactual policies would have had if they were deployed in place of the logging policy.
 In the literature, several estimators have been developed, all with different characteristics and theoretical guarantees.
-Therefore, there is no dominant estimator, and each estimator may be the best one for different OPE problems, depending on the characteristics of the dataset at hand.
-While the selection of the estimator is a crucial choice for an accurate OPE, this problem has been widely overlooked in the literature.
-We propose an automated data-driven OPE estimator selection method based on machine learning.
+Therefore, there is no dominant estimator and each estimator may be the best for different OPE problems, depending on the characteristics of the dataset at hand.
+Although the selection of the estimator is a crucial choice for an accurate OPE, this problem has been widely overlooked in the literature.
+We propose an automated data-driven OPE estimator selection method based on supervised learning.
 In particular, the core idea we propose in this paper is to create several synthetic OPE tasks and use a machine learning model trained to predict the best estimator for those synthetic tasks.
-We empirically show how our method is able to generalize to unseen tasks and make a better estimator selection compared to a baseline method on several real-world datasets, with a computational cost significantly lower than the one of the baseline.
+We empirically show how our method is able to perform a better estimator selection compared to a baseline method on several real-world datasets, with a computational cost significantly lower than the one of the baseline.
 1 INTRODUCTION
 In numerous real-world scenarios, decision-making systems are required to interact with dynamic environments, making choices aimed at optimizing specific objectives.
-These systems are prevalent in domains such as recommender systems [10, 30, 7], information retrieval [22], ad-placement systems [4], or medical treatments [9].
-Such decision-making problems can be viewed through the lens of the Contextual Bandit paradigm [43, 19].
+These systems are prevalent in domains such as recommender systems (Gilotte et al., 2018; Saito and Joachims, 2021; Felicioni et al., 2022), information retrieval (Li et al., 2015), ad-placement systems (Bottou et al., 2013), or medical treatments (Gan et al., 2023).
+Such decision-making problems can be viewed through the lens of the contextual bandit paradigm (Wang et al., 2005; Langford and Zhang, 2007).
 For instance, let us consider the recommendation system domain: the recommender (agent) repeatedly observes user profiles (contexts), and aims to recommend items (actions) to users based on the given context to maximize some notion of reward.
 A critical challenge arises in evaluating the performance of new policies without deploying them in the real world.
-This is especially pertinent in scenarios where deploying a suboptimal policy can have significant negative consequences, such as lost revenue in ad placement or user dissatisfaction in recommender systems.
-Off-Policy Evaluation (OPE) seeks to address this challenge by evaluating the performance of counterfactual policies using data collected from a different, previously deployed policy, often termed the logging policy.
+This is especially pertinent in scenarios where deploying a suboptimal policy can have significant negative consequences.
+Off-Policy Evaluation (OPE) addresses this challenge by evaluating the performance of counterfactual policies using data collected from a different, previously deployed policy, often termed the logging policy.
 The ability to make informed decisions about policy modifications without necessitating real-world deployment is not only cost-effective but also reduces the potential risks associated with untested policies.
 However, the effectiveness of OPE largely hinges on the accuracy of the estimator used to predict the performance of the evaluation policies.
-There is a wide variety of estimators that have been proposed in the literature, each with distinct characteristics and mathematical properties [41].
-As an example, a well-known estimator, called Inverse Propensity Scoring (IPS) [29, 34], mitigates the estimation bias by adjusting observed rewards based on the inverse probability assigned by the logging policy to the logged action.
-In this way, actions that have a low probability under the logging policy receive higher weighting to compensate for their infrequent occurrence in the dataset, and vice versa for actions more likely under the logging policy.
-Despite IPS’s prominence, numerous other estimators exist, each with unique trade-offs between bias and variance.
+There is a wide variety of estimators that have been proposed in the literature, each with distinct characteristics and mathematical properties (Vlassis et al., 2019).
+As an example, a well-known estimator, called Inverse Propensity Scoring (IPS) (Precup et al., 2000; Strehl et al., 2010), provides an unbiased estimation, at the cost of high variance.
+Other estimators also exist, each with unique trade-offs between bias and variance.
 Given the heterogeneity of real-world datasets and the myriad of scenarios in which OPE is applied, no single estimator is universally optimal.
-Further details on related work are reported in Appendix A.
 An estimator may be the best under certain circumstances but perform poorly under others.
@@
 In this paper, we address the Estimator Selection problem using an automated data-driven approach.
-Our main idea consists of exploiting synthetic data.
-We generate several different synthetic OPE tasks by varying the characteristics of each one.
-Various types of estimators are applied to these tasks, and their estimates are compared against the ground-truth value of each evaluation policy, which is known due to the synthetic nature of the data.
-These synthetic OPE tasks are used to train a machine learning model designed to predict estimator’s Mean Squared Errors for the given OPE task.
-This model can select the estimator with the lowest predicted error, effectively addressing the OPE Estimator Selection challenge.
+Our main idea consists of framing this problem as a supervised learning problem.
+In particular, we notice that there are several results suggesting that an estimator’s error is related to the OPE task characteristics.
+Hence, we generate several different synthetic OPE tasks by varying the characteristics of each one, we apply various types of estimators and their estimates are compared against the ground-truth value of each evaluation policy, which is known due to the synthetic nature of the data.
+These synthetic OPE tasks are used to train a supervised model designed to predict estimator’s Mean Squared Error for the given OPE task.
+The supervised model will understand the underlying patterns shared among the various OPE tasks and their relation with the estimator error, and then it can be used to select the estimator with the lowest predicted error, effectively addressing the OPE Estimator Selection challenge.
 We call our Estimator Selection method Automated Off-Policy Estimator Selection (AutoOPE).
-The problem of Estimator Selection has been widely overlooked in the literature, and, to the best of our knowledge, only a very recent paper [40] proposed a method for Estimator Selection, named Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF), which is substantially different from our proposal.
+The problem of Estimator Selection has been widely overlooked in the literature, and, to the best of our knowledge, only a very recent paper (Udagawa et al., 2023) proposed a method for Estimator Selection, named Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF), which is substantially different from our proposal.
 In this paper, we propose a novel data-driven approach called Automated Off-Policy Estimator Selection (AutoOPE) to address the Estimator Selection problem in Off-Policy Evaluation.
@@
 We share the code for our experiments for researchers and practitioners in the field [^1] We anonymously share the code to reproduce our experiments: https://anonymous.4open.science/r/auto-ope-28F3/.
-  • To the best of our knowledge, this is the first work to address the estimator selection problem in OPE using a data-driven approach leveraging only synthetic data, and it is one of the few techniques proposed for Estimator Selection in OPE in general, despite the importance of the problem.
+  • To the best of our knowledge, this is one of the few techniques proposed for Estimator Selection, despite the importance of the problem.
   • We conduct extensive experiments on several datasets that empirically demonstrate the effectiveness of our proposed AutoOPE method in selecting high-performing estimators tailored to the given OPE task, performing better than the baseline with a much reduced computational cost.
-This represents an important contribution towards automated and adaptive Off-Policy Evaluation.
 2 BACKGROUND
@@
 We denote an OPE estimator as $\hat{V}(\pi_{e})$.
-A common metric used to assess the performance of an estimator is the Mean Squared Error (MSE), defined as $MSE(\hat{V}(\pi_{e}))=\mathbb{E}\left[(\hat{V}(\pi_{e})-V(\pi_{e}))^{2}\right]$.
+A common metric used to assess the performance of an estimator is the Mean Squared Error (MSE), defined as $\text{MSE}(\hat{V}(\pi_{e}))=\mathbb{E}\left[(\hat{V}(\pi_{e})-V(\pi_{e}))^{2}\right]$.
 2.3 Estimator Selection for OPE
-The effectiveness of OPE crucially depends on how accurate is the estimator in estimating the policy value of the evaluation policy, usually in terms of how low is its $MSE(\hat{V}(\pi_{e}))$.
+The effectiveness of OPE crucially depends on how accurate is the estimator in estimating the policy value of the evaluation policy, usually in terms of how low is its $\text{MSE}(\hat{V}(\pi_{e}))$.
 For this reason, and since many OPE estimators have been proposed in the literature, selecting the best estimator for a given OPE task, i.e., Estimator Selection, is particularly important.
@@
 Since in OPE we do not have any data collected under the evaluation policy, we cannot compute the ground-truth evaluation policy value $V(\pi_{e})$ and therefore the ground-truth MSE of the estimator.
-A common approach is to approximate it with an estimate, i.e., $\widehat{MSE}(\hat{V}(\pi_{e}))$, and to select the estimator that has the lowest approximate MSE.
+A common approach is to approximate it with an estimate, i.e., $\widehat{\text{MSE}}(\hat{V}(\pi_{e}))$, and to select the estimator that has the lowest approximate MSE.
 2.3.1 Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF)
-To the best of our knowledge, only Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF) [40] has tackled the problem of Estimator Selection in the literature.
+To the best of our knowledge, only Policy-Adaptive Estimator Selection via Importance Fitting (PAS-IF) (Udagawa et al., 2023) has tackled the problem of Estimator Selection in the literature.
 To accurately estimate the MSE of an OPE estimator, the key idea of PAS-IF is to determine a sampling rule $\rho_{\theta}:\mathcal{X}\times\mathcal{A}\rightarrow(0,1)$, that governs a subsampling procedure to divide the logging data $\mathcal{D}_{b}$ into two pseudo datasets: a pseudo logging dataset $\tilde{\mathcal{D}}_{b}$ and a pseudo evaluation dataset $\tilde{\mathcal{D}}_{e}$.
@@
 After that, each estimator MSE can be approximated in this way:
-$\widehat{MSE}_{\mbox{PAS-IF}}(\hat{V}(\pi_{e}))\coloneqq(\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})-\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b}))^{2}\ .$
+$\widehat{\text{MSE}}_{\mbox{PAS-IF}}(\hat{V}(\pi_{e}))\coloneqq(\hat{V}_{on}(\pi_{e};\mathcal{\tilde{D}}_{e})-\hat{V}(\pi_{e};\mathcal{\tilde{D}}_{b}))^{2}\ .$
 One limitation of PAS-IF is that $\rho_{\theta}(x_{t},a_{t})$ is different for every OPE task, hence requiring to re-train a separate PAS-IF model for each OPE task.
@@
 3.1 Theoretical Motivation
-AutoOPE is theoretically justified by many findings that show how the performance of an OPE estimator is highly dependent on the characteristics of the OPE task at hand [42, 32].
-For instance, [24, 25, 26] theoretically analyze Inverse Propensity Scoring (IPS) [29, 34], one of the most used OPE estimators and the basis for many others.
-Metelli et al.
-[24] show that the IPS error is strictly related to the (exponentiated) Rényi divergence between the logging and evaluation policies [^2] The Rényi divergence is a divergence between two probability distributions, and it is defined as $d_{r}(\pi_{e}||\pi_{b})=\log\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\ a\sim\pi_{b}}\left[(\pi_{e}(a|x)/\pi_{b}(a|x))^{2}\right]$.
-and provide a finite-sample bound on the absolute estimation error of IPS under standard assumptions:
-$\left\lvert\hat{V}_{\text{IPS}}(\pi_{e})-V(\pi_{e})\right\rvert\leq\mathcal{O}\left(\sqrt{\frac{\exp(d_{r}(\pi_{e}||\pi_{b}))}{n_{b}}}\right),$
-later also proving that this bound is tight [26].
-This theoretical finding, along with many others, shows how the performance of an OPE estimator is strictly related to the characteristics of the OPE task, such as the divergence between the evaluation and the logging policy.
-In other words, we assume that:
-$MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})=f^{*}(g(\hat{V},\mathcal{T}_{\text{ope}}))\ ,$
-which means that there is a function $f^{*}$ that express the relationship between the real error of an estimator on a given OPE task $MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})$ and the features $g(\hat{V},\mathcal{T}_{\text{ope}})$ of the estimator and the task.
-For this reason, we propose our zero-shot meta-learning approach [8], which we frame as a supervised learning problem.
-Let us assume that we have a large variety of tasks at our disposal $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$, a set of estimators $\mathcal{V}$, and the feature extraction function $g$.
-Also, let us assume that we have access to the ground-truth MSE.
-If we have access to these, we could try to find the function that best predicts the MSE for all tasks and estimators, as a standard supervised problem:
-$\hat{f}=\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)\ ,$
-where $\mathcal{L}$ is a suitable loss.
-Once we found $\hat{f}$, we can apply it in a zero-shot way whenever we need to perform estimator selection with a new task $\mathcal{T}_{\text{ope}}^{(\text{new})}\notin\mathcal{M}$.
-Indeed, with $\hat{f}$ we can now estimate the MSE for a given task in this way:
-$\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right)\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})).$
+AutoOPE is motivated by many findings that show how the performance of an OPE estimator is highly dependent on the characteristics of the OPE task at hand Voloshin et al.
+(2021); Saito et al.
+(2021b).
+To make one example, Metelli et al.
+(2018) show that the IPS error is strictly related to the (exponentiated) Rényi divergence (with $\alpha=2$) between the logging and evaluation policies [^2] The Rényi divergence is a family of divergences between two probability distributions with a parameter $\alpha$.
+For $\alpha=2$, it is defined as $d_{r}(\pi_{e}||\pi_{b})=\log\operatorname*{\mathbb{E}}_{x}\operatorname*{\mathbb{E}}_{\ a\sim\pi_{b}}\left[(\pi_{e}(a|x)/\pi_{b}(a|x))^{2}\right]$.
+and provides a finite sample bound on the absolute estimation error of IPS under standard assumptions: $\left\lvert\hat{V}_{\text{IPS}}(\pi_{e})-V(\pi_{e})\right\rvert\leq\mathcal{O}\left(\sqrt{\frac{\exp(d_{r}(\pi_{e}||\pi_{b}))}{n_{b}}}\right).$ This theoretical finding is just one example among many others that shows how the performance of an OPE estimator is related to the characteristics of the OPE task.
+In other words, we make the following assumption:
+Assumption 1.
+Consider a fixed set of estimators $\mathcal{V}$.
+There exists a function $f^{*}\in\mathcal{F}$ that expresses the relationship between the real error of any estimator on a given OPE task $MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})$ and the features $g(\hat{V},\mathcal{T}_{\text{ope}})$ of the estimator and the task:
+$\text{MSE}(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}})=f^{*}(g(\hat{V},\mathcal{T}_{\text{ope}})),\quad\forall\hat{V}\in\mathcal{V}.$
+Let us assume that we have a variety of tasks at our disposal $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$, a fixed set of estimators $\mathcal{V}$, and the feature extraction function $g$.
+With these, we can frame a supervised learning problem and apply empirical risk minimization to find the function that best predicts the MSE for all tasks and estimators:
+$\begin{split}\hat{f}=&\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{\begin{subarray}{c}\hat{V}\\ \mathcal{T}_{\text{ope}}\end{subarray}}\mathcal{L}\Bigl{(}\text{MSE}(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\Bigr{)},\end{split}$
+where $\mathcal{L}$ is a suitable loss, $\mathcal{F}$ is the hypothesis class, and $N=|\mathcal{V}||\mathcal{M}|$.
+Now, we can prove that the empirical risk minimizer $\hat{f}$ is accurate with high probability.
+To provide an intuition, we consider a finite hypothesis class, but one can extend this finding to infinite hypothesis classes via standard learning theory tools (e.g., VC-dimension, Rademacher complexity) (Shalev-Shwartz and Ben-David, 2014).
+Also, we consider a subgaussian loss, which is less restrictive than only considering bounded loss functions (Germain et al., 2016).
+Assumption 2 (Sub-Gaussian Loss).
+The loss function $\mathcal{L}$ is subgaussian with parameter $\sigma^{2}$.
+Specifically, for all $y,\hat{y}$, the random variable $\mathcal{L}(y,\hat{y})-\mathbb{E}[\mathcal{L}(y,\hat{y})]$ satisfies:
+$\mathbb{E}\left[e^{\lambda(\mathcal{L}(y,\hat{y})-\mathbb{E}[\mathcal{L}(y,\hat{y})])}\right]\leq e^{\frac{\lambda^{2}\sigma^{2}}{2}}\quad\forall\lambda\in\mathbb{R}.$
+Assumption 3 (Finite hypothesis class).
+The hypothesis class $\mathcal{F}$ is finite: $|\mathcal{F}|<\infty$.
+Theorem 1.
+Under Assumptions 1, 2, and 3, suppose we have $N=|\mathcal{V}||\mathcal{M}|$ independent samples $\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}=g(\hat{V}_{i},\mathcal{T}_{i})$ and $y_{i}=MSE(\hat{V}_{i}(\pi_{e});\mathcal{T}_{i})$, drawn uniformly over the estimators $\hat{V}\in\mathcal{V}$ and tasks $\mathcal{T}\in\mathcal{M}$.
+Let $\hat{f}$ be the empirical risk minimizer over $\mathcal{F}$:
+$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
+Then, for any $\delta>0$, with probability at least $1-\delta$, the expected loss $L(\hat{f})$ satisfies:
+$L(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}},$
+where $\sigma^{2}$ is the subgaussian parameter from Assumption 2, and the expected loss $L(f)$ is defined as
+$L(f)=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f(x)\right)\right].$
+Proof.
+The proof derives from mapping the estimator selection problem to a supervised learning one, and then applying standard learning theory tools, such as Hoeffding’s inequality and the union bound.
+See Appendix B for all the deferred proofs.
+∎
+Theorem 1 shows that, with high probability, the empirical risk minimizer $\hat{f}$ has a low error, and the error term that decreases with $N$.
+Since the number of estimator is fixed, the error term decreases the number of tasks $|\mathcal{M}|$ contained in the dataset.
+This theoretical finding is intuitive, and it shows that $\hat{f}$ is trained to understand the patterns of the given tasks and it learns more if provided with more tasks.
+Once we have found $\hat{f}$, we can apply it in a zero-shot way whenever we need to perform estimator selection with a new task $\mathcal{T}_{\text{ope}}^{(\text{new})}\notin\mathcal{M}$.
+With $\hat{f}$ we can now estimate the MSE for a given task in this way:
+$\widehat{MSE}_{\mbox{AutoOPE}}(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})})).$
 Hence, we just need to select the estimator that minimizes the MSE estimated by AutoOPE:
-$\hat{V}^{*}=\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}\right).$
+$\hat{V}^{*}=\arg\min_{\hat{V}\in\mathcal{V}}\widehat{MSE}_{\mbox{AutoOPE}}(\hat{V},\mathcal{T}_{\text{ope}}^{(\text{new})}).$
 3.2 Practical Implementation
-However, this approach is feasible if we have access to several OPE tasks, a feature extraction function, and the ground-truth MSE.
-For this reason, our approach extensively rely on synthetic data.
+In the previous section, we provided a theoretical intuition for our approach.
+In this section, we describe how to implement the proposed solution in practice.
+In particular, this approach needs several OPE tasks, a feature extraction function, and the ground-truth MSE.
+For this reason, we rely on synthetic data.
 Specifically:
 Dataset
-We generate a meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ composed of several different synthetic OPE tasks.
+We generate a dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ composed of several different synthetic OPE tasks.
 In this way, we have access to the ground-truth MSE.
@@
 We indeed generate 250,000 synthetic OPE tasks.
-Furthermore, to favor generalization of our model, we create OPE tasks with different numerical characteristics.
-For more details on the synthetic data generation, we refer to Appendix C.
+Furthermore, to favor generalization of our model, we create OPE tasks with different distributions.
+For more details on the synthetic data generation, we refer to Appendix D.
 Features
@@
 (iii) OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, or uses importance sampling.
-See Appendix E for their full description.
+See Appendix F for their full description.
 Estimators
-The set $\mathcal{V}$ of candidate estimators include several model-free, model-based and hybrid estimators, as done in [40]. All the estimators used are described in Appendix B.
+The set $\mathcal{V}$ of candidate estimators include several model-free, model-based and hybrid estimators, as done in Udagawa et al. (2023). All the estimators used are described in Appendix C.
 Model
-As a supervised model $\hat{f}$, we use a Random Forest regressor [5] because of its high-quality predictions on tabular data and its speed, both in training and inference.
-For more details on training and optimization, we refer to Appendix D.
-We recap the AutoOPE method in Algorithms 1 (training of the model) and 2 (zero-shot estimator selection).
+As a supervised model $\hat{f}$, we use a Random Forest regressor (Breiman, 2001) because of its high-quality predictions on tabular data and its speed, both in training and inference.
+For more details on training and optimization, we refer to Appendix E.
 Algorithm 1 AutoOPE training
@@
 0:  MSE estimator $\widehat{MSE}_{\mbox{AutoOPE}}$
-1:  Generate meta-dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ (Appendix C)
+1:  Generate dataset $\mathcal{M}\coloneqq\{\mathcal{T}_{\text{ope}}^{(1)},\mathcal{T}_{\text{ope}}^{(2)},\dots\}$ (Appendix D)
 2:  $\mathcal{L}\leftarrow$ Random Forest regression loss
-3:  $\hat{f}\leftarrow\arg\min_{f}\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M}}\sum_{\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)$ (Train the random forest)
+3:  $\hat{f}\leftarrow\arg\min_{f}$ $\sum_{\mathcal{T}_{\text{ope}}\in\mathcal{M},\hat{V}\in\mathcal{V}}\mathcal{L}\left(MSE(\hat{V}(\pi_{e});\mathcal{T}_{\text{ope}}),f(g(\hat{V},\mathcal{T}_{\text{ope}}))\right)$ (Train the random forest)
 4:  For any $\hat{V},\mathcal{T}_{\text{ope}}$, define $\widehat{MSE}_{\mbox{AutoOPE}}\left(\hat{V},\mathcal{T}_{\text{ope}}\right)\coloneqq\hat{f}(g(\hat{V},\mathcal{T}_{\text{ope}}))$
@@
 2:  return  $\hat{V}^{*}$
+We recap the AutoOPE method in Algorithms 1 (training of the model) and 2 (zero-shot estimator selection).
 The AutoOPE approach has several advantages.
-First of all, it is data-driven: AutoOPE tries to learn the shared patterns in several different synthetic OPE tasks with the objective to generalize to new unseen tasks.
-For this, we use a Random Forest, which is a powerful and fast model for tabular data.
-Furthermore, it is theory-inspired: it stems from theoretical findings to create a data-driven model that takes in input features such as divergences which are proven to be related to the estimator’s error, but also features that have no theoretical guarantee at the moment.
-Thanks to this, we can also provide insights that can shed light on future theoretical findings by analyzing the features that had a significant impact on the model’s performance.
+First of all, it is data-driven: AutoOPE tries to learn the shared patterns in several different OPE tasks.
+Furthermore, it is theory-inspired: we create a data-driven model inspired by the fact that there are some features of the OPE tasks which are theoretically proven to be related to the estimator’s error.
 Finally, it is zero-shot.
 This means that a practitioner can avoid training for each new estimator selection task, leading to a much faster computational time.
-A possible downside of our approach is that we trained AutoOPE on large meta-dataset $\mathcal{M}$ that is synthetic and created a priori, which may not lead to generalization to particular real-world OPE tasks.
-However, in Section 4, we show that AutoOPE generalizes well to unseen real-world OPE tasks with an extensive experimental evaluation on a large number of real-world OPE tasks.
+A possible downside of our approach is that we trained AutoOPE on a synthetic data set $\mathcal{M}$ created a priori, which may not lead to generalization to particular real-world OPE tasks.
+In particular, Theorem 1, provides the guarantee that our method generalizes on the same distribution of tasks.
+This implicitly assumes that, at test time, the OPE tasks come from the same distribution of the synthetic training tasks.
+However, this assumption might not hold.
+For this reason, we relax this assumption in the following:
+Theorem 2.
+Let $P_{S}$ and $P_{T}$ be the distributions of synthetic and real (test) tasks respectively.
+Let $L_{S}(f)$ and $L_{T}(f)$ be the expected loss of $f$ with respect to $P_{S}$ and $P_{T}$ respectively.
+Under the same assumptions of Theorem 1, for any $\delta>0$, with probability at least $1-\delta$, we can bound $L_{T}$ of the empirical risk minimizer $\hat{f}$ as follows:
+$L_{T}(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}$
+where $d$ is the chi-squared divergence, defined as $d(P_{S},P_{T})=\operatorname*{\mathbb{E}}_{P_{S}}[((P_{T}(x)-P_{S}(x))/P_{S}(x))^{2}]$, and $K$ is a constant.
+Also this finding is intuitive: we have that our approach learns to generalize on the test OPE tasks as long as it is trained on a large number of synthetic tasks and the divergence $d(P_{S},P_{T})$ is small.
+Notice that, the real OPE task distribution is unknown, and so is $d(P_{S},P_{T})$.
+In order to find whether $d(P_{S},P_{T})$ is small in practice or not, in Section 4, we test AutoOPE on unseen real-world OPE tasks with an extensive experimental evaluation [^3] We test on almost 50 different real-world experimental configurations, while prior work is usually tested on 2 or 3.
+on a large number of real-world OPE tasks, obtained from different real datasets.
 4 EXPERIMENTS
-In this section, we describe the experimental analysis of our proposed AutoOPE and show it is competitive with the PAS-IF baseline on several real-world tasks, as well as exhibiting a lower computational cost.
+In this section, we describe the experimental analysis of our proposed AutoOPE and show its strong performance compared the PAS-IF baseline on several real-world tasks, while exhibiting a lower computational cost.
 4.1 Experimental Settings
@@
 [^4] We estimate that applying the same Bayesian Search used for AutoOPE would require at least a month of high-end GPU usage.
-For this reason we rely on the ad-hoc strategy presented in [40] which tunes the regularization parameter $\lambda$ for each new OPE task via grid search.
+For this reason we rely on the ad-hoc strategy presented in (Udagawa et al., 2023) which tunes the regularization parameter $\lambda$ for each new OPE task via grid search.
 AutoOPE
-We train the Random Forest model of AutoOPE on the synthetic meta-dataset $\mathcal{M}$ and we optimize its hyperparameters with a Bayesian search.
-For more details, we refer to Appendix D.
+We train the Random Forest model of AutoOPE on the synthetic dataset $\mathcal{M}$ and we optimize its hyperparameters with a Bayesian search.
+For more details, we refer to Appendix E.
 In this Section, we only use the trained Random Forest in inference, in a zero-shot way (see Algorithm 2) on new unseen tasks.
 Evaluation Metrics for OPE Selection
-To evaluate AutoOPE and PAS-IF on the Estimator Selection task we use two metrics: Relative Regret and Spearman’s Rank Correlation Coefficient, as in [40], leveraging the fact that in all experiments we have access to the ground-truth MSE of each estimator.
-Further details are in Appendix H.
+To evaluate AutoOPE and PAS-IF on the Estimator Selection task we use two metrics: Relative Regret and Spearman’s Rank Correlation Coefficient, as in (Udagawa et al., 2023), leveraging the fact that in all experiments we have access to the ground-truth MSE of each estimator.
+Further details are in Appendix J.
 […]
 […]
-Spearman’s Rank Correlation Coefficient [33] measures the similarity between two rankings, one based on the MSE predicted by the Estimator Selection method and the other based on the ground-truth MSE.
+Spearman’s Rank Correlation Coefficient (Spearman, 1961) measures the similarity between two rankings, one based on the MSE predicted by the Estimator Selection method and the other based on the ground-truth MSE.
 4.2 Results
-Here we present the results we obtained on several public real-world datasets.
 4.2.1 Open Bandit Dataset (OBD)
-In this real-world experiment, we use the Open Bandit Dataset (OBD) [31], a publicly available dataset originated from a large-scale fashion e-commerce platform.
+In this real-world experiment, we use the Open Bandit Dataset (OBD) (Saito et al., 2021a), a publicly available dataset originated from a large-scale fashion e-commerce platform.
 The dataset encompasses three campaigns: ‘ALL’, ‘MEN’, and ‘WOMEN’.
@@
 To further strengthen our empirical evaluation of the generalization capabilities of AutoOPE, in this section we present 40 other real-world experiments conducted on 8 different datasets.
-In particular, we took 8 datasets from the UCI repository [1], as typically done in the OPE literature [2, 35, 6, 32].
+In particular, we took 8 datasets from the UCI repository (Asuncion and Newman, 2007), as typically done in the OPE literature (Beygelzimer and Langford, 2009a; Su et al., 2020a; Dudík et al., 2015; Saito et al., 2021b).
 We report an overview of these datasets in Table 2.
@@
 CB Data
-To adapt these datasets for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features as context features and the class labels as actions, while the reward is 1 if the current action is the correct class for the given context, 0 vice versa.
+To adapt these datasets for CB we perform the standard supervised to bandit conversion (Beygelzimer and Langford, 2009a; Joachims et al., 2018), reinterpreting the input features as context features and the class labels as actions, while the reward is 1 if the current action is the correct class for the given context, 0 vice versa.
 Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies.
@@
 As expected, AutoOPE is in the order of 10 times faster compared to PAS-IF, even though PAS-IF performs estimator selection on a GPU, while AutoOPE does it on a CPU.
-Further details are reported in Appendix G.
+Further details are reported in Appendix I.
 4.4 Additional Experiments
-Finally, we conducted other additional experiments on more datasets, various ablation studies, as well as an analysis of the most important features of the AutoOPE model.
-Feature Importance
-A useful benefit of using Random Forest models is that we can compute the feature importance.
-For this analysis we use the Mean Decrease in Impurity (MDI), meaning that features with higher importance scores are more influential in the model decision-making process.
 Figure 3 shows a selection of features for the Random Forest model.
-The full plot is in Figure 4 of Appendix E.
-To assess the importance of each feature group, let us focus on the 15 most and least important features in the full plot of Figure 4 in Appendix E.
-The 11 policy-independent features appear to be of limited importance and correspond to 4 of the 15 least important features.
-Only 1 policy-dependent feature belongs to the first 15 most important features, which is the 4th moment of the logged rewards and is the second most important feature.
-Policy-dependent features, which are 24, are generally useful for the model, with 10 among the 15 most important features and 6 out of 24 among the 15 least important.
-The KL divergence is the most important feature, which suggests further theoretical study may be warranted.
-The 8 OPE estimator features are similarly spread, with 4 among the most important features and 5 among the least important ones.
-In particular, among the most important OPE estimator features there are features that characterize the type of estimator (if it uses a reward model or importance sampling or both).
-The insights gained from this feature importance analysis is an important contribution.
-In particular, it can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
-Additional Experiments and Ablation studies
-We conducted additional experiments on estimator selection on synthetic datasets and on the CIFAR dataset [18].
-Also, we conducted different ablation studies, investigating the impact of scaling the size of the meta-dataset, the diversity of the meta-dataset, and the features used.
-We describe all these additional experiments in Appendix F.
+The full plot is in Figure 4, Appendix F, together with its discussion.
+The insights gained from this feature importance analysis can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
+Also, we conducted additional experiments on estimator selection on synthetic datasets and on the CIFAR dataset (Krizhevsky et al., 2009) reported in Appendix G.
+Finally, we conducted different ablation studies, investigating the impact of scaling the size of the synthetic dataset, the diversity of the dataset, and the features used.
+We describe all these ablation experiments in Appendix H.
 5 CONCLUSIONS
@@
 Through several experiments, we empirically showed how our method is able to consistently make a better estimator selection when compared to the baseline method on several real-world datasets, outperforming the baseline method by a large margin with a significantly lower computational cost.
-For future work, a natural direction to follow would be to try different machine learning models for AutoOPE, in particular for scenarios where one may want further guarantees on the Estimator Selection process.
-Furthermore, one could try to theoretically analyze the relationship between the features of the AutoOPE meta-dataset and the performance of the various estimators, which was unveiled by the feature importance of the random forest model.
+For future work, a natural direction to follow would be to try different machine learning models for AutoOPE.
+Furthermore, one could try to theoretically analyze the relationship between the features of the AutoOPE dataset and the performance of the various estimators, which was unveiled by the feature importance of the random forest model.
 Finally, our approach focused on the estimator selection problem in the contextual bandit domain.
-Our method could be used as a stepping stone towards improving OPE estimator selection in the more general Reinforcement Learning scenario.
+Our method could be used as a stepping stone towards OPE estimator selection in the more general Reinforcement Learning scenario.
 Appendix A RELATED WORK
-In the realm of Contextual Bandit Off-Policy Evaluation, there exists a rich and diverse landscape of published research aimed at assessing the performance of policies without actually deploying them, avoiding the hazard of selecting action under not valuable or deleterious policies which, depending on the scenario, could be costly and dangerous [10, 21, 30, 17].
-Indeed, this area of study has brought contributions to various application domains, including recommender systems [22], advertising [4], and precision medicine [23], reflecting its flexibility in modeling different situations.
-In this context, AutoOPE emerges as a novel approach, offering a unique solution based on a black-box model that leverages synthetic data and meta-learning to tackle the Estimator Selection problem.
+In the realm of Contextual Bandit Off-Policy Evaluation, there exists a rich and diverse landscape of published research aimed at assessing the performance of policies without actually deploying them, avoiding the hazard of selecting action under not valuable or deleterious policies which, depending on the scenario, could be costly and dangerous (Gilotte et al., 2018; Levine et al., 2020; Saito and Joachims, 2021; Kiyohara et al., 2021).
+Indeed, this area of study has brought contributions to various application domains, including recommender systems (Li et al., 2015), advertising (Bottou et al., 2013), and precision medicine (Lu et al., 2021), reflecting its flexibility in modeling different situations.
+In this context, AutoOPE emerges as a novel approach, offering a unique solution based on a supervised learning model that leverages synthetic data and supervised learning to tackle the Estimator Selection problem.
 To appreciate the significance of AutoOPE, it is essential to understand the broader context of Contextual Bandit OPE research.
 Several pioneering works have laid the foundation for evaluating counterfactual policies in the context of Contextual Bandit.
-Notable contributions include studies that have introduced fundamental approaches such as the Direct Method (DM) [3], Inverse Propensity Scoring (IPS) [29, 34], and Doubly Robust (DR) [38, 12, 6], which serve as the cornerstone for OPE research, since they define the three main classes of OPE estimators: model-based, model-free and hybrid.
+Notable contributions include studies that have introduced fundamental approaches such as the Direct Method (DM) (Beygelzimer and Langford, 2009b), Inverse Propensity Scoring (IPS) (Precup et al., 2000; Strehl et al., 2010), and Doubly Robust (DR) (Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015), which serve as the cornerstone for OPE research, since they define the three main classes of OPE estimators: model-based, model-free and hybrid.
 Direct Method (DM) is rooted in machine learning, employing algorithms to estimate policy performance by regressing rewards.
@@
 However, its performance still depends on the specific task so it is not necessarily the best option.
-A multitude of OPE estimators, has been designed over time as extensions or combinations of DM, IPS and DR [37, 14, 35, 26, 38, 44]. These efforts collectively contribute to advance OPE methodologies.
-However, many estimators incorporate hyperparameters to balance the bias-variance trade-off given the Contextual Bandit environment such as [35, 26, 38, 12, 6, 44].
-As reported by [32, 31] the hyperparameter tuning strongly affects the prediction quality.
-To address the issue of hyperparameter optimization [36] and [39] proposed SLOPE and its improved version SLOPE++, based on Lepski’s principle [20], offering improved results compared to naïve tuning procedures.
+A multitude of OPE estimators, has been designed over time as extensions or combinations of DM, IPS and DR (Swaminathan and Joachims, 2015; Kallus and Uehara, 2019; Su et al., 2020a; Metelli et al., 2021; Thomas and Brunskill, 2016; Wang et al., 2017).
+These efforts collectively contribute to advance OPE methodologies.
+However, many estimators incorporate hyperparameters to balance the bias-variance trade-off given the Contextual Bandit environment such as (Su et al., 2020a; Metelli et al., 2021; Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015; Wang et al., 2017).
+As reported by (Saito et al., 2021b, a) the hyperparameter tuning strongly affects the prediction quality.
+To address the issue of hyperparameter optimization (Su et al., 2020b) and (Tucker and Lee, 2021) proposed SLOPE and its improved version SLOPE++, based on Lepski’s principle (Lepski and Spokoiny, 1997), offering improved results compared to naïve tuning procedures.
 Lepski’s principle allows to define an upper-bound for the Mean Squared Error without delving into the bias estimation, that is the main reason why the task of MSE estimation is challenging.
 SLOPE dominates all the other methods designed for hyperparameter tuning, being the state of the art technique for this task.
-Other existing methods like MAGIC [38], or other built-in methods of proposed estimators [35, 44] have tackled hyperparameter selection using MSE estimators, but they offer weaker guarantees compared to SLOPE.
+Other existing methods like MAGIC (Thomas and Brunskill, 2016), or other built-in methods of proposed estimators (Su et al., 2020a; Wang et al., 2017) have tackled hyperparameter selection using MSE estimators, but they offer weaker guarantees compared to SLOPE.
 The Estimator Selection problem has emerged as a relatively unexplored but pivotal aspect of OPE research.
-Empirical studies by [42] and [32] have shown the fundamental role of Estimator Selection on OPE accuracy, highlighting its importance in practical applications.
-In this context, the recent PAS-IF (Policy-Adaptive Estimator Selection via Importance Fitting) method [40] emerges as the only baseline for Estimator Selection.
+Empirical studies by (Voloshin et al., 2021) and (Saito et al., 2021b) have shown the fundamental role of Estimator Selection on OPE accuracy, highlighting its importance in practical applications.
+In this context, the recent PAS-IF (Policy-Adaptive Estimator Selection via Importance Fitting) method (Udagawa et al., 2023) emerges as the only baseline for Estimator Selection.
 PAS-IF uses a neural network to learn a subsampling rule that is used to generate two pseudo-datasets starting from the available logging data.
@@
 Once the two pseudo datasets have been generated, PAS-IF allows to estimate the candidate estimators’ Mean Squared Error based on the empirical average of rewards in these pseudo datasets.
-Our proposed AutoOPE method, on the other hand, competes with PAS-IF in performance by introducing a novel data-driven approach that leverages meta-learning with synthetic data and a black-box model.
+Our proposed AutoOPE method, on the other hand, competes with PAS-IF in performance by introducing a novel data-driven approach that leverages learning from synthetic data with a supervised model.
 AutoOPE offers a versatile approach, allowing the comparison of different estimator classes and the selection of the most accurate one.
 This innovative approach empirically promises to significantly enhance the accuracy of OPE, reducing an important gap in the Contextual Bandit Off-Policy Evaluation field.
-Appendix B Off-policy Estimators Description
+Appendix B DEFERRED PROOFS
+Theorem 1.
+Under Assumptions 1, 2, and 3, suppose we have $N=|\mathcal{V}||\mathcal{M}|$ independent samples $\{(x_{i},y_{i})\}_{i=1}^{N}$, where each $x_{i}=g(\hat{V}_{i},\mathcal{T}_{i})$ and $y_{i}=\text{MSE}(\hat{V}_{i}(\pi_{e});\mathcal{T}_{i})$, drawn uniformly over the estimators $\hat{V}\in\mathcal{V}$ and tasks $\mathcal{T}\in\mathcal{M}$.
+Let $\hat{f}$ be the empirical risk minimizer over the finite hypothesis class $\mathcal{F}$:
+$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
+Then, for any $\delta>0$, with probability at least $1-\delta$, the expected loss $L(\hat{f})$ satisfies:
+$L(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}},$
+where $\sigma^{2}$ is the subgaussian parameter from Assumption 2, and the expected loss $L(f)$ is defined as
+$L(f)=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f(x)\right)\right].$
+Proof.
+Under Assumption 1, there exists a function $f^{*}\in\mathcal{F}$ such that for all $(\hat{V},\mathcal{T})$,
+$y=\text{MSE}(\hat{V}(\pi_{e});\mathcal{T})=f^{*}(x)=f^{*}(g(\hat{V},\mathcal{T})).$
+This implies that the expected loss of $f^{*}$ is zero:
+$L(f^{*})=\mathbb{E}_{(x,y)}\left[\mathcal{L}\left(y,f^{*}(x)\right)\right]=0.$
+Let $\Delta(f)=L_{\mathcal{M}}(f)-L(f)$ denote the deviation between the empirical and expected losses, where
+$L_{\mathcal{M}}(f)=\frac{1}{N}\sum_{i=1}^{N}\mathcal{L}\left(y_{i},f(x_{i})\right).$
+By Assumption 2, for any fixed $f\in\mathcal{F}$, the random variable $\mathcal{L}\left(y_{i},f(x_{i})\right)$ is subgaussian with parameter $\sigma^{2}$.
+Using Hoeffding’s inequality for subgaussian variables, we have for any $\epsilon>0$:
+$\mathbb{P}\left(|\Delta(f)|\geq\epsilon\right)\leq 2\exp\left(-\frac{N\epsilon^{2}}{2\sigma^{2}}\right).$
+Since $\mathcal{F}$ is finite, applying the union bound over all hypotheses:
+$\mathbb{P}\left(\sup_{f\in\mathcal{F}}|\Delta(f)|\geq\epsilon\right)\leq 2|\mathcal{F}|\exp\left(-\frac{N\epsilon^{2}}{2\sigma^{2}}\right).$
+Setting the right-hand side equal to $\delta$ and solving for $\epsilon$, we get:
+$\epsilon=\sigma\sqrt{\frac{2\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}.$
+On the event where $\sup_{f\in\mathcal{F}}|\Delta(f)|\leq\epsilon$, the following holds for the empirical risk minimizer $\hat{f}$:
+$L(\hat{f})=L_{\mathcal{M}}(\hat{f})-\Delta(\hat{f})\leq L_{\mathcal{M}}(\hat{f})+\epsilon.$
+Since $L_{\mathcal{M}}(\hat{f})\leq L_{\mathcal{M}}(f^{*})$ by definition of $\hat{f}$, and $L_{\mathcal{M}}(f^{*})=L(f^{*})+\Delta(f^{*})=0+\Delta(f^{*})\leq\epsilon$, we have:
+$L(\hat{f})\leq L_{\mathcal{M}}(\hat{f})+\epsilon\leq L_{\mathcal{M}}(f^{*})+\epsilon\leq\epsilon+\epsilon=2\epsilon.$
+Substituting $\epsilon$ yields the desired bound:
+$L(\hat{f})\leq 2\sigma\sqrt{\frac{2\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}=\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}.$
+Theorem 2.
+Let $P_{S}$ and $P_{T}$ be the distribution of synthetic and real test tasks respectively.
+Let $L_{S}(f)$ and $L_{T}(f)$ be the expected loss of $f$ with respect to $P_{S}$ and $P_{T}$ respectively.
+Under the same assumptions of Theorem 1, for any $\delta>0$, with probability at least $1-\delta$, we can bound $L_{T}$ of the empirical risk minimizer $\hat{f}$ as follows:
+$L_{T}(\hat{f})\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}$
+where $d$ is a divergence defined as $d(P_{S},P_{T})=\operatorname*{\mathbb{E}}_{P_{S}}[((P_{T}(x)-P_{S}(x))/P_{S}(x))^{2}]$, and $K$ is a constant.
+Proof.
+For any $f\in\mathcal{F}$:
+$L_{T}(f)$ $=L_{S}(f)+L_{T}(f)-L_{S}(f)$ (2)
+$\leq L_{S}(f)+\left\lvert L_{T}(f)-L_{S}(f)\right\rvert$ (apply absolute value)
+[…]
+[…]
+[…]
+[…]
+$=L_{S}(f)+K\sqrt{d(P_{S},P_{T})}$
+where $d$ is the chi-squared divergence, defined as:
+[…]
+If we consider the empirical risk minimizer $\hat{f}$ on $\mathcal{M}\overset{\mathrm{iid}}{\sim}P_{S}$ defined as:
+$\hat{f}=\arg\min_{f\in\mathcal{F}}\frac{1}{N}\sum_{(x_{i},y_{i})\in\mathcal{M}}\mathcal{L}\left(y_{i},f(x_{i})\right),$
+and we apply Theorem 1, we have that:
+$L_{T}(\hat{f})$ $\leq L_{S}(\hat{f})+K\sqrt{d(P_{S},P_{T})}$ (applying Eq. 2)
+$\leq\sqrt{\frac{8\sigma^{2}\log\left(\dfrac{2|\mathcal{F}|}{\delta}\right)}{N}}+K\sqrt{d(P_{S},P_{T})}\ ,$ (applying Theorem 1)
+with probability at least $1-\delta$.
+Appendix C OFF-POLICY ESTIMATORS DESCRIPTION
 This Appendix contains a brief characterization of all the OPE estimators, $\mathcal{V}$, used in this paper.
 Note that the total number of OPE candidate estimators is 21 which includes 3 model-free estimators and 6 requiring a reward predictors, of which three were included.
-B.1 Model-Based Estimators
+C.1 Model-Based Estimators
 We include one estimator based exclusively on a reward model.
 Direct Method (DM)
-The DM estimator [3] employs a supervised machine learning model $\hat{q}$ to estimate the mean reward $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$.
+The DM estimator (Beygelzimer and Langford, 2009b) employs a supervised machine learning model $\hat{q}$ to estimate the mean reward $\mathbb{E}_{p(r_{t}|x_{t},a_{t})}\left[r_{t}\right]$.
 The policy value prediction is computed in this way:
@@
 It can be a good choice when the complexity of the task is not too high.
-B.2 Model-Free Estimators
+C.2 Model-Free Estimators
 We include three estimators that do not rely on a reward predictor, but instead are built on the statistical concept of importance sampling.
 Inverse Propensity Scoring (IPS)
-The IPS estimator [29, 34] employs pure importance sampling to re-weight the average of rewards, accounting for the distribution shift between $\pi_{b}$ and $\pi_{e}$, since the average of rewards are collected under the logging policy, while the goal is to evaluate $\pi_{e}$.
+The IPS estimator (Precup et al., 2000; Strehl et al., 2010) employs pure importance sampling to re-weight the average of rewards, accounting for the distribution shift between $\pi_{b}$ and $\pi_{e}$, since the average of rewards are collected under the logging policy, while the goal is to evaluate $\pi_{e}$.
 The policy value prediction is computed in this way:
@@
 Self-Normalized Inverse Propensity Scoring (SNIPS)
-The SNIPS estimator [37, 14] tries to stabilize the IPS estimator normalizing the prediction by the empirical mean of the importance weights. The policy value prediction is computed in this way:
+The SNIPS estimator (Swaminathan and Joachims, 2015; Kallus and Uehara, 2019) tries to stabilize the IPS estimator normalizing the prediction by the empirical mean of the importance weights.
+The policy value prediction is computed in this way:
 $\hat{V}_{SNIPS}(\pi_{e};\mathcal{T}_{\text{ope}})\coloneqq\frac{\sum_{t}^{n_{b}}w(x_{t},a_{t})r_{t}}{\sum_{t}^{n_{b}}w(x_{t},a_{t})}\ \ .$
@@
 Subgaussian Importance Sampling Inverse Propensity Scoring (IPS-$\lambda$)
-The IPS-$\lambda$ estimator [26] shrink the importance ratio smoothly, with a sub-gaussian parameterized by a value $\lambda$ to reduce the IPS variance.
+The IPS-$\lambda$ estimator (Metelli et al., 2021) shrink the importance ratio smoothly, with a sub-gaussian parameterized by a value $\lambda$ to reduce the IPS variance.
 The policy value prediction is computed in this way:
@@
 The $\lambda,\ \gamma$ hyperparameters balance the bias-variance trade-off of the estimator.
-B.3 Hybrid Estimators
+C.3 Hybrid Estimators
 We include five hybrid estimators, that leverage both a reward predictor and importance sampling.
 Doubly Robust (DR)
-The DR estimator [38, 12, 6] employs a reward model $\hat{q}$ like DM, but it improves the reward prediction by $\hat{q}$ using a correction term based on importance sampling.
+The DR estimator (Thomas and Brunskill, 2016; Jiang and Li, 2016; Dudík et al., 2015) employs a reward model $\hat{q}$ like DM, but it improves the reward prediction by $\hat{q}$ using a correction term based on importance sampling.
 The policy value prediction is computed in this way:
@@
 Self-Normalized Doubly Robust (SNDR)
-The SNDR estimator [38, 37, 14] is a variant of DR that shrinks the weights using the same normalization factor used in SNIPS. The policy value prediction is computed in this way:
+The SNDR estimator (Thomas and Brunskill, 2016; Swaminathan and Joachims, 2015; Kallus and Uehara, 2019) is a variant of DR that shrinks the weights using the same normalization factor used in SNIPS.
+The policy value prediction is computed in this way:
 $\hat{V}_{SNDR}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q})\coloneqq\sum_{t=1}^{n_{b}}\left(\frac{w(x_{t},a_{t})}{\sum_{j=1}^{n}w(x_{j},a_{j})}(r_{t}-q(x_{t},a_{t}))+\frac{1}{n_{b}}\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right)\ \ .$
@@
 Subgaussian Importance Sampling Doubly Robust (DR-$\lambda$)
-The DR-$\lambda$ estimator [26] is a variant of DR that shrinks the weights using the same smooth subgaussian by IPS-$\lambda$. The policy value prediction is computed in this way:
+The DR-$\lambda$ estimator (Metelli et al., 2021) is a variant of DR that shrinks the weights using the same smooth subgaussian by IPS-$\lambda$. The policy value prediction is computed in this way:
 $\hat{V}_{DR-\lambda}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q},\lambda,\gamma)\coloneqq\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(w_{\lambda,\gamma}(x_{t},a_{t})(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right)\ \ .$
@@
 Doubly Robust with Optimistic Shrinkage (DRos)
-The DRos estimator [35] is a variant of DR that reduces variance with an importance weight derived by minimizing the sharp bound of estimator MSE. The policy value prediction is computed in this way:
+The DRos estimator (Su et al., 2020a) is a variant of DR that reduces variance with an importance weight derived by minimizing the sharp bound of estimator MSE.
+The policy value prediction is computed in this way:
 $\hat{V}_{DRos}(\pi_{e};\mathcal{T}_{\text{ope}},\hat{q},\lambda)\coloneqq\frac{1}{n_{b}}\sum_{t=1}^{n_{b}}\left(\frac{\lambda w(x_{t},a_{t})}{w(x_{t},a_{t})^{2}+\lambda}(r_{t}-q(x_{t},a_{t}))+\sum_{a\in\mathcal{A}}\pi_{e}(a|x_{t})\hat{q}(x_{t},a)\right),$
@@
 Switch
-The Switch estimator [44] is a variant of DR that switches between DM and DR, using DM when the importance ratio is below a threshold specified by the value of $\lambda$, and DR when $w(x_{t},a_{t})$ is above the threshold.
+The Switch estimator (Wang et al., 2017) is a variant of DR that switches between DM and DR, using DM when the importance ratio is below a threshold specified by the value of $\lambda$, and DR when $w(x_{t},a_{t})$ is above the threshold.
 The policy value prediction is computed in this way:
@@
 where the $\lambda\geq 0$ hyperparameter balances the bias-variance trade-off of the estimator.
-Appendix C Meta-Dataset
-C.1 Characteristics
-As stated in Section 3, AutoOPE method is based on a zero-shot meta-learning approach framed as a supervised learning problem.
+Appendix D SYNTHETIC DATASET
+D.1 Characteristics
+As stated in Section 3, AutoOPE method is based on a zero-shot approach framed as a supervised learning problem.
 In particular, the goal of this supervised learning problem is to predict, given an estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$ as input, the MSE of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$.
 Hence, to train a Random Forest regressor to predict this MSE, we need a finite set of possible estimators $\mathcal{V}$ such that $\hat{V}_{m}\in\mathcal{V}$ defined a-priori, and a large set of OPE tasks $\mathcal{M}$.
-The set $\mathcal{M}$ is a collection of synthetically generated OPE tasks, obtained from a data generator function $G$, which we describe in further detail in Section C.2.
+The set $\mathcal{M}$ is a collection of synthetically generated OPE tasks, obtained from a data generator function $G$, which we describe in further detail in Section D.2.
 The training dataset for the Random Forest is obtained extracting suitable input features from the estimator $\hat{V}_{m}$ and an OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, that the Random Forest can work with.
-This is done using a feature extraction function $g$.
+This is done using a feature extraction function $g$ (described in Appendix F.).
 Then, the target is the ground-truth MSE of $\hat{V}_{m}$ when applied to $\mathcal{T}_{\text{ope}}^{(j)}$.
-The training dataset can be expressed as: $\{g(\mathcal{V}\times\mathcal{M}),\ MSE(\mathcal{V}\times\mathcal{M})\}$
-The meta-learning dataset $\mathcal{M}$ contains $250,000$ OPE tasks, each associated to all the candidate estimators in $\mathcal{V}$ that are $21$, for a total of $5.2$ million samples.
+The training dataset can be expressed as: $\{\mathrm{input:}\ g(\mathcal{V}\times\mathcal{M}),\ \mathrm{target:}\ \text{MSE}(\mathcal{V}\times\mathcal{M})\}$
+The meta-dataset $\mathcal{M}$ contains $250,000$ OPE tasks, and for each task $\mathcal{T}_{\text{ope}}^{(j)}$, $n_{gen}=10$ realizations $\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}}$ are produced, for a total of $2,500,000$ OPE tasks.
+Each task is associated to all the candidate estimators in $\mathcal{V}$ that are $21$, for a total of $52.5$ million samples.
 Note that we removed the few trivial OPE tasks which had all rewards equal to 0 or 1.
 OPE Estimators
-The set $\mathcal{V}$ of candidate estimators includes several model-free, model-based and hybrid estimators, as done in [40].
-Estimators hyperparameters, when present, were selected using the SLOPE method [36].
-All the estimators used are listed in Table 3, and a brief explanation of each of them is reported in appendix B.
-Each of the six estimators with a reward model is employed in conjunction with three reward predictors: Random Forest Classifier and Logistic Regression from the Scikit-Learn [^4] Scikit-Learn available with 3-Clause BSD license: https://github.com/scikit-learn/scikit-learn library [28], and LGBM Classifier from the Python package of LightGBM [^5] LightGBM available with MIT license: https://github.com/microsoft/LightGBM/tree/master [16].
-The reward models were fitted using a 3-fold cross-fitting procedure to reduce possible overfitting [27].
+The set $\mathcal{V}$ of candidate estimators includes several model-free, model-based and hybrid estimators, as done in (Udagawa et al., 2023).
+Estimators hyperparameters, when present, were selected using the SLOPE method (Su et al., 2020b).
+All the estimators used are listed in Table 3, and a brief explanation of each of them is reported in appendix C.
+Each of the six estimators with a reward model is employed in conjunction with three reward predictors: Random Forest Classifier and Logistic Regression from the Scikit-Learn [^5] Scikit-Learn available with 3-Clause BSD license: https://github.com/scikit-learn/scikit-learn library (Pedregosa et al., 2011), and LGBM Classifier from the Python package of LightGBM [^6] LightGBM available with MIT license: https://github.com/microsoft/LightGBM/tree/master (Ke et al., 2017).
+The reward models were fitted using a 3-fold cross-fitting procedure to reduce possible overfitting (Narita et al., 2021).
 Type OPE Estimator
@@
 Switch DR (SwitchDR)
-Table 3: List of the OPE estimators used in this paper. See Appendix B for their description.
-Features
-We compute 43 features for each OPE task that can be categorized into three groups:
-  • Policies-independent features, that describe structural properties of the CB scenario and are unrelated from the specific $\pi_{b},\ \pi_{e}$, e.g., the number of rounds, the reward mean and variance, the number of actions.
-  • Policies-dependent features, such as the dissimilarity between the two policies $\pi_{b},\ \pi_{e}$, measured with different geometrical and statistical distances.
-  • OPE estimator features, that refer to specific properties of the OPE estimator used in that task, e.g., if the estimator is self-normalized, uses importance sampling, employs a reward model, and if so, which one.
-The comprehensive list of features, together with their mathematical formulations, is reported in Appendix E.
+Table 3: List of the OPE estimators used in this paper. See Appendix C for their description.
 Target
-The target variable is the MSE between the prediction of an estimator $\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(j)})\ \text{with}\ \hat{V}_{m}\in\mathcal{V}$ and the on-policy estimated ground-truth policy value $V(\pi_{e})$.
-Indeed, together with each synthetic OPE task $\mathcal{T}_{\text{ope}}^{(j)}$, we also generate a ground-truth dataset, denoted as $\mathcal{D}_{gt}^{(j)}\coloneqq\left\{(x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{gt}}$, that is a large CB dataset collected under the evaluation policy $\pi_{e}$.
-Realizations
-In order to improve the robustness of the data, both the features and MSE are computed as the average of $n_{gen}$ realizations $\mathcal{T}_{ope}^{(j)}(s)$ of each $\mathcal{T}_{ope}^{(j)}$.
-The features and the target MSE are computed for each candidate estimator and averaged across all realizations to obtain the final values used in the meta-learning dataset.
-C.2 Generation
+The target variable is the MSE between the prediction of an estimator $\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(j)}(s))\ \text{with}\ \hat{V}_{m}\in\mathcal{V}$ and the on-policy estimated ground-truth policy value $V(\pi_{e})$.
+Indeed, together with each synthetic OPE task $\mathcal{T}_{\text{ope}}^{(j)}(s)$, we also generate a ground-truth dataset, denoted as $\mathcal{D}_{gt}^{(j)}\coloneqq\left\{(x_{t},a_{t},r_{t},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{gt}}$, that is a large CB dataset collected under the evaluation policy $\pi_{e}$, which is used to compute the ground-truth MSE, by averaging the collected rewards (see Appendix J for further details).
+D.2 Generation
 Data Generator
-We can formalize the generation of the pair $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})$ as the output of a function $G$ which serves as synthetic data generator, so $(\mathcal{T}_{\text{ope}}^{(j)},\mathcal{D}_{gt}^{(j)})=G(p_{1}^{(j)},\dots,p_{P}^{(j)})$, with $(p_{1}^{(j)},\dots,p_{P}^{(j)})$ drawn from $\Phi=(\phi_{1}(\cdot),\dots,\phi_{P}(\cdot))$.
+Synthetizing the previous section, we can say that it is needed a pair $(\mathcal{T}_{\text{ope}}^{(j)}(s),\mathcal{D}_{gt}^{(j)})$ and one estimator $\hat{V}_{m}$, to generate one row of the dataset on which the Random Forest will be trained.
+We can formalize the generation of the pair ($(\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}},\mathcal{D}_{gt}^{(j)})$ as the output of a function $G$ which serves as synthetic data generator, so $(\mathcal{T}_{\text{ope}}^{(j)}(s),\mathcal{D}_{gt}^{(j)})=G(p_{1}^{(j)},\dots,p_{P}^{(j)})$, with the parameters $(p_{1}^{(j)},\dots,p_{P}^{(j)})$ drawn from $\Phi=(\phi_{1}(\cdot),\dots,\phi_{P}(\cdot))$.
 For each new OPE task, we sample a value for each parameter $p_{i}$ from the respective distribution $\phi_{i}$, and generate the datasets pair.
@@
 Among the parameters $\left\{p_{i}\right\}_{i}^{P}$ that change the OPE task characteristics we include: the evaluation and logging policies, the number of possible actions, the number of rounds in the logging data, the context dimension, the reward distribution.
-Below we provide the complete list of these parameters, along with their distributions $\left\{\phi_{i}\right\}_{i}^{P}$, providing insight into the characteristics considered for OPE tasks:
+Below we provide the complete list of these parameters, along with their distributions $\left\{\phi_{i}\right\}_{i}^{P}$, providing insights into the characteristics considered for OPE tasks:
   • $n_{a}$: a positive integer, representing the number of possible actions available in the considered Contextual Bandit scenario.
@@
 Apart from these randomly sampled parameters, the synthetic data generator $G$ has additional parameters that are fixed and independent from the current generated OPE task. These include:
-  • $n_{data}$: a positive integer determining the number of synthetic OPE task data ($\mathcal{T}_{\text{ope}}$) to generate. In total, we generated $n_{data}=250,000$ different OPE task.
   • $n_{gen}$: a positive integer determining the number of realizations of each $\mathcal{T}_{\text{ope}}$. In our generation process, $n_{gen}=10$.
@@
 OPE Task Generation
-Introduced the parameters used in the generation process, the procedure for obtaining the two datasets $(\mathcal{T}_{\text{ope}},\mathcal{D}_{gt})$ can be formally described.
-The procedure is outlined by Algorithm 3. Internally to the function Generate, called in the Algorithm 3:
+Introduced the parameters used in the generation process, the procedure for obtaining the pair ($(\{\mathcal{T}_{\text{ope}}^{(j)}(s)\}_{s}^{n_{gen}},\mathcal{D}_{gt}^{(j)})$ can be formally described.
+The procedure is outlined by Algorithm 3, which internally calls the function Generate and the function GenerateOPE, explained in the following:
+Generate
   1. (1)
@@
     Then actions $a_{t}$ are sampled for each context following policy $\pi_{e}$, and the relative rewards $r_{t}$ for each pair $(x_{t},a_{t})$ are sampled from a Bernoulli distribution with a mean of $q(x_{t},a_{t})$.
-Afterwards, within the function GenerateOPE, called in the Algorithm 3:
+  5. (5)
+    The data produced are sufficient to formally define a logging dataset to compute the on-policy value estimate as rewards average.
+GenerateOPE
   1. (1)
@@
     The correspondent reward $r_{t}$ is produced and collected, sampling from the same Bernoulli distribution we introduced before.
-These data are sufficient to formally define $\mathcal{T}_{\text{ope}}$ as in Equation 1.
-The procedure to generate $\mathcal{T}_{\text{ope}}$ is repeated $n_{gen}$ times, obtaining $n_{gen}$ realizations of the $\mathcal{T}_{\text{ope}}$ dataset.
+  5. (5)
+    The data produced are sufficient to formally define $\mathcal{T}_{\text{ope}}^{(j)}(s)$ as in Equation 1.
 Algorithm 3 OPE Data Generator $G$
-0:  number of OPE task generations $n_{data}$, number of realizations $n_{gen}$ of each OPE dataset, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$ for the evaluation dataset, space of generation parameters $\Phi$
-0:  Collection of OPE datasets and evaluation datasets $\left\{\{\mathcal{T}^{(j)}_{ope}(s)\}_{s}^{n_{gen}},\ \mathcal{D}^{(j)}_{gt}\right\}_{j=1}^{n_{data}}$
-1:  for $j=1,\ldots,n_{data}$ do
-2:     $(n_{a},n_{b},d_{x},q,\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}})\sim\Phi$
-3:     $\mathcal{A}\leftarrow\left\{0,\dots,n_{a}-1\right\}$
-4:     $\mathcal{D}^{(j)}_{gt}\leftarrow\mathrm{Generate}(\mathcal{A},d_{x},\beta_{e},f_{\pi_{e}},q,n_{gt})$
-5:     for $s=1,\ldots,n_{gen}$ do
-6:        $\mathcal{T}^{(j)}_{ope}(s)\leftarrow\mathrm{GenerateOPE}(\mathcal{A},d_{x},\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
-7:     end for
-8:  end for
-Appendix D Supervised Model Training
+0:  number of realizations $n_{gen}$ of each OPE dataset, a big number of rounds $n_{gt}$ to sample under $\pi_{e}$ for the evaluation dataset, space of generation parameters $\Phi$
+0:  Collection of OPE datasets and evaluation datasets $\{\mathcal{T}^{(j)}_{ope}(s)\}_{s}^{n_{gen}},\ \mathcal{D}^{(j)}_{gt}\}_{j=1}^{n_{data}}$
+1:  $(n_{a},n_{b},d_{x},q,\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}})\sim\Phi$
+2:  $\mathcal{A}\leftarrow\left\{0,\dots,n_{a}-1\right\}$
+3:  $\mathcal{D}^{(j)}_{gt}\leftarrow\mathrm{Generate}(\mathcal{A},d_{x},\beta_{e},f_{\pi_{e}},q,n_{gt})$
+4:  for $s=1,\ldots,n_{gen}$ do
+5:     $\mathcal{T}^{(j)}_{ope}(s)\leftarrow\mathrm{GenerateOPE}(\mathcal{A},d_{x},\beta_{b},\beta_{e},f_{\pi_{b}},f_{\pi_{e}},q,n_{b},s)$
+6:  end for
+Appendix E SUPERVISED MODEL TRAINING
 Supervised Model
 As a supervised model, we select the Random Forest regressor provided by the scikit-learn package [^8] https://scikit-learn.org/stable/modules/generated/sklearn.ensemble.RandomForestRegressor.html.
+The learning pipeline comprehend also some preprocessing steps and transformations:
+  • Categorical features are one-hot encoded.
+  • Numerical features are clipped to a maximum value of $10^{10}$.
+  • The function $log(1+x)$ is applied to positive features with a skewness higher than 1 (and to the target variable).
+  • Each numerical feature is scaled to the interval $(0,1)$ if the feature is positive and to $(-1,1)$ if can assume negative values. The scaling is applied also the target variable.
 Data Split
-In order to train and evaluate the AutoOPE supervised model, we split the meta-learning dataset in 80% for optimization and training, and 20% for testing, so that the testing data never contains OPE tasks the supervised model was trained on.
+In order to train and evaluate the AutoOPE supervised model, we split the meta-dataset into three parts, 60% for training, 20% for validation, and 20% for testing.
 Hyperparameter Optimization
 AutoOPE uses a Random Forest as a supervised model.
-In order to select the best hyperparameters we rely on a Bayesian Search optimizing the 5-fold cross-validation Regret between the predicted MSE of the OPE estimator selected by AutoOPE, $\hat{V}_{\hat{m}}$, and the best estimator according to the MSE on the ground-truth dataset.
+In order to select the best hyperparameters we rely on a Bayesian Search optimizing the Regret between the ground-truth MSE of the OPE estimator selected by AutoOPE, $\hat{V}_{\hat{m}}$, and the best estimator according to the MSE on the ground-truth dataset.
 We define the Regret as:
-$\text{Regret}\coloneqq MSE(\hat{V}_{\hat{m}}(\pi_{e}))-MSE(\hat{V}_{m^{*}}(\pi_{e}))\ ,$ (4)
-This means that at each step of the Bayesian Search, the training dataset is split into five folds, and a Random Forest regression model (with fixed hyperparameters) is fitted for each of the possible five subsets of four data folds, resulting in five models in total.
-Each model is then evaluated on the remaining out-of-bag fold in terms of Regret.
-The Regret values are averaged over the different five models, and the 5-fold cross-validation metric is obtained, that is the function optimized by the Bayesian Search.
+$\text{Regret}\coloneqq\text{MSE}(\hat{V}_{\hat{m}}(\pi_{e}))-\text{MSE}(\hat{V}_{m^{*}}(\pi_{e}))\ ,$ (5)
+This means that at each step of the Bayesian Search, a Random Forest regression model (with fixed hyperparameters) is fitted on the training set, and then is evaluated on the validation set in terms of Regret.
+This Regret is the objective function that the Bayesian Search tries to optimize.
+Finally, the Random Forest is trained on the union of the training and validation set with the best hyperparameters found by the Bayesian Search, and is evaluated on the test set to asses the performance.
 The Bayesian Search is implemented on top of the Scikit-Optimize [^9] Scikit-Optimize available with 3-Clause BSD license: https://github.com/scikit-optimize/scikit-optimize/tree/master library, and involves Bayesian optimization using Gaussian Processes to approximate the objective function, making the assumption that function points are distributed as a multivariate Gaussian.
-The Bayesian Optimization is executed on the training data, with 100 calls made to optimize the hyperparameters.
+For our model, 50 Bayesian Search iterations are performed.
 Type Hyperparameter Range Prior
-Integer n_estimator [200] -
+Integer n_estimator (50, 500) -
 max_depth (1, 100) uniform
@@
 Table 4: Hyperparameters Search Space for AutoOPE Random Forest model Hyperparameters Tuning via Bayesian Search
-Appendix E Features Design
-This Appendix mathematically describes the input features for the Random Forest in AutoOPE. They are:
+Appendix F FEATURE DESIGN
+This Appendix mathematically describes the input features computed by the feature extractor $g$. They are:
   • Policy-Independent features:
@@
     • Neyman chi-squared distance between $\pi_{b},\pi_{e}$, which we included because it is equivalent (up to a constant) to the exponentiated Rényi divergence (shown to be an important data characteristic by Metelli et al.
-[24]):
+(2018)):
       $\text{Neyman}(\pi_{b}||\pi_{e})=\exp(d_{r}(\pi_{e}||\pi_{b}))-1\ \ .$
@@
     • Whether the estimator uses a Logistic Regression as reward model or not.
+Feature Importance
 In Figure 4 is reported the complete Feature Importance plot, measured with Mean Decrease in Impurity (MDI).
+The 10 policy-independent features appear to be of limited importance and correspond to 5 of the 15 least important features.
+Only 2 policy-independent features belong to the first 15 most important features.
+Policy-dependent features, which are 24, are generally useful for the model, with 9 among the 15 most important features and 6 out of 24 among the 15 least important.
+The KL divergence is the most important feature, which confirms what previous studies in theory have found (Chatterjee and Diaconis, 2018).
+The 9 OPE estimator features are similarly spread, with 4 among the most important features and 4 among the least important ones.
+The insights gained from this feature importance analysis is another contribution.
+In particular, it can shed light on future theoretical findings, lead by analyzing the features that had a significant impact on the model’s performance.
 Figure 4: Plot of the Feature Importance measured with Mean Decrease in Impurity (MDI). The most important features are at the top.
-Appendix F Additional Experiments and Ablation Studies
-F.1 Scaling Experiment
-We performed a scaling experiment of AutoOPE by varying the size of the meta-learning dataset $\mathcal{M}$.
+Appendix G ADDITIONAL EXPERIMENTS
+G.1 Additional Experiments on CIFAR-10
+The CIFAR-10 dataset (Krizhevsky et al., 2009) is a widely-used benchmark for image classification tasks, with mutually exclusive labeling.
+CB Data
+To adapt the CIFAR-10 dataset for CB we perform the standard supervised to bandit conversion (Beygelzimer and Langford, 2009a; Joachims et al., 2018), reinterpreting the input features (i.e., images pixels scaled in $(0,1)$) as context features and the class labels as actions.
+Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies.
+In order to generate the policies, the second data split is used to train two Logistic Regression classifiers.
+Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0.
+To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using using parameters $\alpha_{b}$, $\alpha_{e}$:
+$\pi_{b}(a|x)\coloneqq\alpha_{b}\cdot\pi_{det,b}(a|x)+(1-\alpha_{b})\cdot\pi_{u}(a|x)\ \ \ ,$
+$\pi_{e}(a|x)\coloneqq\alpha_{e}\cdot\pi_{det,e}(a|x)+(1-\alpha_{e})\cdot\pi_{u}(a|x)\ \ \ ,$
+[…]
+Results
+Figure 5: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection of experiments on CIFAR-10 dataset.
+Shaded areas correspond to 95% confidence intervals.
+In order to be able to compute the variance, we perform 20 bootstrap samples of the logging data and we perform the experimental tests on each of them.
+The results are shown in Figure 5.
+AutoOPE consistently outperforms PAS-IF, demonstrating lower Relative Regret and better Spearman’s Rank Correlation Coefficient.
+Notably, AutoOPE exhibits a much lower variance in particular for the Relative Regret, while PAS-IF exhibits very high variance.
+These results are consistent with the experiments showed in Section 4 and confirm that AutoOPE is able to generalize to unseen OPE tasks with a different distribution compared to the one used for its training, in a zero-shot scenario.
+G.2 Additional Synthetic Experiments
+Figure 6: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 1.
+Shaded areas correspond to 95% confidence intervals.
+Figure 7: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 2.
+Shaded areas correspond to 95% confidence intervals.
+In this section we describe all the experiments on synthetic data, following the experimental pipeline of (Udagawa et al., 2023).
+CB Data
+The synthetic experiments are based on two different logging policies, Logging 1 and Logging 2.
+For each of them 21 OPE tasks with different evaluation policies are defined via a parameter $\beta_{e}$.
+The synthetic logging data is generated using the following procedure and distributions (same procedure used in (Udagawa et al., 2023)):
+  • Context vectors are generated as 10-dimensional vectors, sampled from a multivariate normal distribution with zero mean and identity covariance matrix.
+  • A total of 10 possible actions is considered.
+  • For each context, we compute the expected reward $\mathbb{E}\left[r(x_{t},a_{t})\right]$, for all available actions.
+These expected rewards are determined using a synthetic reward function from the ‘dataset’ module of OBP, specifically the ‘logistic_reward_function’, described by Equation 3.
+The logging policies $\pi_{b}(a_{t}|x_{t})$ in these experiments are defined as a combination of two policies, each represented as $\pi_{i}(a_{t}|x_{t})=\textit{softmax}_{a_{t}}(\beta_{i}\cdot q(x_{t},a_{t}))$, with $i\in\left\{1,2\right\}$: the first half of rounds is generated under $\pi_{1}$, and the second half under $\pi_{2}$.
+In total, $n_{b,i}=1,000$ contexts are drawn for each policy $\pi_{i}$, resulting in a total number of logging rounds $n_{b}=2000$.
+We compute the logging distribution $\pi_{b}(a_{t}|x_{t})$ of all actions, for each context $x_{t}$ sampled, and afterwards, we randomly sample an action $a_{t}$ for each $x_{t}$, according to this distribution.
+The actual rewards for these pairs $(x_{t},a_{t})$ are sampled from a Bernoulli distribution with a mean of $q(x_{t},a_{t})$.
+This process yields a series of context-action-reward triplets $\left\{(x_{t},a_{t},r_{t})\right\}_{t}^{n_{b}}$.
+These triplets, together with the action distribution $\pi_{b}$ computed for all the $x_{t}$, constitute the logging dataset $\mathcal{D}_{b}=\left\{(x_{t},a_{t},r_{t},\left\{\pi_{b}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{b}}$.
+The evaluation policy $\pi_{e}$ is described using the same expression as the logging policies, namely $\pi_{e}(a_{t}|x_{t})=\text{softmax}_{a_{t}}(\beta_{e}\cdot q(x_{t},a_{t}))$, and the counterfactual action distribution is computed for each context $x_{t}$ in $\mathcal{D}_{b}$.
+It can be noticed that, with this evaluation policy definition, the $\pi_{e}$ is uniquely determined by the parameter $\beta_{e}$.
+At the end, the complete OPE dataset generated for a synthetic experiment is described as $\mathcal{T}_{\text{ope}}\coloneqq\left\{\mathcal{D}_{b},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$.
+As already said, we divide the synthetic experiments into two main settings: one named Logging 1, where we set $\beta_{1}=2$ and $\beta_{2}=-2$, and the other called Logging 2, where $\beta_{1}=3$ and $\beta_{2}=7$, in line with the experimental analysis performed by Udagawa et al.
+(2023).
+By varying $\beta_{e}$ values in a range from -10 to 10 for both Logging 1 and 2, we consider 21 different OPE experiments in each setting.
+To provide precise performance assessments, in all the 21 experiments of both Logging 1 and 2 we generate $n_{data}=100$ realizations of the logging dataset, as done in the procedure outlined in (Udagawa et al., 2023).
+The performance metrics are subsequently averaged over all realizations to yield more robust performance estimates with their confidence intervals.
+Results
+The results obtained from Logging 1 are reported in Figure 7 while those of Logging 2 in Figure 7.
+Since these synthetic experiments are the same of (Udagawa et al., 2023), we use their optimal values for the other hyperparameters such as the network structure, regularization, learning rate.
+In both cases our proposed AutoOPE outperforms PAS-IF across almost all evaluation policies in terms of Spearman’s rank correlation coefficient.
+In terms of Relative Regret, AutoOPE performs similar to PAS-IF in the first half settings of the Logging 1 experiments.
+However, for high values of $\beta_{e}$, PAS-IF slightly outperform AutoOPE.
+In the Logging 2 set of experiments, we see that for all the $\beta_{e}\geq 0$ AutoOPE outperforms PAS-IF, but PAS-IF performs better in the settings with $\beta_{e}<0$.
+The fact that AutoOPE exhibits better Spearman’s Rank Correlation indicates that, while both methods predict similar values for the best estimator, AutoOPE is more consistent in ranking all of them.
+Appendix H ABLATION STUDIES
+In this appendix, we report some ablation studies.
+Due to the specific settings of these experiments, several new model optimizations need to be executed.
+Given the limited time and resources, we consider a smaller version of the meta-dataset $\mathcal{M}$ to produce results within a reasonable timeframe.
+For the same reason, we fix the number of trees in the Random Forest to 200, which we find to be a reasonable trade-off between power and manageability.
+In the original dataset, each OPE task generated has 10 realizations, with a row of feature values for each realization.
+In this smaller version of the meta-dataset, however, we consider only one row per OPE task, obtained by averaging the features of the 10 realizations.
+These experiments aim to demonstrate that limiting certain characteristics of the meta-dataset’s OPE tasks, such as the number of actions or the KL-divergence of the policies, can affect model performance.
+Importantly, the conclusions obtained from the following experiments remain valid, independently of the meta-dataset size.
+H.1 Scaling Experiment
+We performed a scaling experiment of AutoOPE by varying the size of the dataset $\mathcal{M}$.
 We take 20% of the OPE tasks in $\mathcal{M}$ as the test set.
 To investigate the scaling behavior, we train and optimize AutoOPE on different sub-samples of the remaining 80% of the OPE tasks in $\mathcal{M}$.
-The results are reported in Figure 5 and show that the loss (i.e., the Regret) of AutoOPE follows a scaling behavior, which is particularly popular in the recent Deep Learning literature [15, 11].
-By increasing the training data, we observe a decrease in the Regret approximately following a power law in the form: $L=A+\frac{B}{D^{\alpha}}$, where $L$ is the loss measured with the Regret, $D$ is the meta-learning dataset size, and $A$, $B$, and $\alpha$ are scalar coefficients.
-While we selected the largest meta-learning dataset for the experimental analysis of this paper, the results in Figure 5 show that the dataset size can be adjusted based on the trade-off between computational resources and model performance for the scenario of interest.
-Figure 5: Scaling experiment by varying the meta-learning dataset size.
-F.2 Additional Ablation Studies
+The results are reported in Figure 8 and show that the loss (i.e., the Regret) of AutoOPE follows a scaling behavior, which is particularly popular in the recent Deep Learning literature (Kaplan et al., 2020; Hoffmann et al., 2022).
+By increasing the training data, we observe a decrease in the Regret approximately following a power law in the form: $L=A+\frac{B}{D^{\alpha}}$, where $L$ is the loss measured with the Regret, $D$ is the meta-dataset size, and $A$, $B$, and $\alpha$ are scalar coefficients.
+While we selected the largest meta-dataset for the experimental analysis of this paper, the results in Figure 8 show that the dataset size can be adjusted based on the trade-off between computational resources and model performance for the scenario of interest.
+Figure 8: Scaling experiment by varying the meta-dataset size.
+H.2 Additional Ablation Studies
 Features
@@
   • estimators-dependent.
-We train 3 models (one for each group) in the same way outlined in Appendix D, and we test them on the same unseen test set used for the scaling experiment. We show the results in Table 5.
+We train 3 models (one for each group) in the same way outlined in Appendix E, and we test them on the same unseen test set used for the scaling experiment. We show the results in Table 5.
 $\begin{array}[]{l|l|l}\hline\cr\text{Features subset}&\text{Rel.
@@
 From these findings, we can see how a more diverse generation method is beneficial and improves performance compared to a generation method that generates tasks with a low number of actions or with similar logging and evaluation policies.
-F.3 Additional Experiments on CIFAR-10
-The CIFAR-10 dataset [18] is a widely-used benchmark for image classification tasks, with mutually exclusive labeling.
-CB Data
-To adapt the CIFAR-10 dataset for CB we perform the standard supervised to bandit conversion [2, 13], reinterpreting the input features (i.e., images pixels scaled in $(0,1)$) as context features and the class labels as actions.
-Furthermore, we partition the dataset into two subsets, the first one is used as logging data $\mathcal{D}_{b}$ for the OPE task, while the second is used to generate both logging and evaluation policies.
-In order to generate the policies, the second data split is used to train two Logistic Regression classifiers.
-Each classifier defines a deterministic policy $\pi_{det}$ because, given a context, it predicts a specific class label (action) with a probability of 1, while all other labels having a probability of 0.
-To produce the final stochastic policies we blend the deterministic policy with a uniform random policy $\pi_{u}(a|x)$ using using parameters $\alpha_{b}$, $\alpha_{e}$:
-$\pi_{b}(a|x)\coloneqq\alpha_{b}\cdot\pi_{det,b}(a|x)+(1-\alpha_{b})\cdot\pi_{u}(a|x)\ \ \ ,$
-$\pi_{e}(a|x)\coloneqq\alpha_{e}\cdot\pi_{det,e}(a|x)+(1-\alpha_{e})\cdot\pi_{u}(a|x)\ \ \ ,$
-[…]
-Results
-Figure 6: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection of experiments on CIFAR-10 dataset.
-Shaded areas correspond to 95% confidence intervals.
-In order to be able to compute the variance, we perform 20 bootstrap samples of the logging data and we perform the experimental tests on each of them.
-The results are shown in Figure 6.
-AutoOPE consistently outperforms PAS-IF, demonstrating lower Relative Regret and better Spearman’s Rank Correlation Coefficient.
-Notably, AutoOPE exhibits a much lower variance in particular for the Relative Regret, while PAS-IF exhibits very high variance.
-These results are consistent with the experiments showed in Section 4 and confirm that AutoOPE is able to generalize to unseen OPE tasks with a different distribution compared to the one used for its training, in a zero-shot scenario.
-F.4 Additional Synthetic Experiments
-Figure 7: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 1.
-Shaded areas correspond to 95% confidence intervals.
-Figure 8: Relative Regret (lower is better) and Spearman’s Rank Correlation Coefficient (higher is better) for Off-Policy Estimator Selection in synthetic experiments with Logging 2.
-Shaded areas correspond to 95% confidence intervals.
-In this section we describe all the experiments on synthetic data, following the experimental pipeline of [40].
-CB Data
-The synthetic experiments are based on two different logging policies, Logging 1 and Logging 2.
-For each of them 21 OPE tasks with different evaluation policies are defined via a parameter $\beta_{e}$.
-The synthetic logging data is generated using the following procedure and distributions (same procedure used in [40]):
-  • Context vectors are generated as 10-dimensional vectors, sampled from a multivariate normal distribution with zero mean and identity covariance matrix.
-  • A total of 10 possible actions is considered.
-  • For each context, we compute the expected reward $\mathbb{E}\left[r(x_{t},a_{t})\right]$, for all available actions.
-These expected rewards are determined using a synthetic reward function from the ‘dataset’ module of OBP, specifically the ‘logistic_reward_function’, described by Equation 2.
-The logging policies $\pi_{b}(a_{t}|x_{t})$ in these experiments are defined as a combination of two policies, each represented as $\pi_{i}(a_{t}|x_{t})=\textit{softmax}_{a_{t}}(\beta_{i}\cdot q(x_{t},a_{t}))$, with $i\in\left\{1,2\right\}$: the first half of rounds is generated under $\pi_{1}$, and the second half under $\pi_{2}$.
-In total, $n_{b,i}=1,000$ contexts are drawn for each policy $\pi_{i}$, resulting in a total number of logging rounds $n_{b}=2000$.
-We compute the logging distribution $\pi_{b}(a_{t}|x_{t})$ of all actions, for each context $x_{t}$ sampled, and afterwards, we randomly sample an action $a_{t}$ for each $x_{t}$, according to this distribution.
-The actual rewards for these pairs $(x_{t},a_{t})$ are sampled from a Bernoulli distribution with a mean of $q(x_{t},a_{t})$.
-This process yields a series of context-action-reward triplets $\left\{(x_{t},a_{t},r_{t})\right\}_{t}^{n_{b}}$.
-These triplets, together with the action distribution $\pi_{b}$ computed for all the $x_{t}$, constitute the logging dataset $\mathcal{D}_{b}=\left\{(x_{t},a_{t},r_{t},\left\{\pi_{b}(a|x_{t})\right\}_{a\in\mathcal{A}})\right\}_{t}^{n_{b}}$.
-The evaluation policy $\pi_{e}$ is described using the same expression as the logging policies, namely $\pi_{e}(a_{t}|x_{t})=\text{softmax}_{a_{t}}(\beta_{e}\cdot q(x_{t},a_{t}))$, and the counterfactual action distribution is computed for each context $x_{t}$ in $\mathcal{D}_{b}$.
-It can be noticed that, with this evaluation policy definition, the $\pi_{e}$ is uniquely determined by the parameter $\beta_{e}$.
-At the end, the complete OPE dataset generated for a synthetic experiment is described as $\mathcal{T}_{\text{ope}}\coloneqq\left\{\mathcal{D}_{b},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$.
-As already said, we divide the synthetic experiments into two main settings: one named Logging 1, where we set $\beta_{1}=2$ and $\beta_{2}=-2$, and the other called Logging 2, where $\beta_{1}=3$ and $\beta_{2}=7$, in line with the experimental analysis performed by Udagawa et al.
-[40].
-By varying $\beta_{e}$ values in a range from -10 to 10 for both Logging 1 and 2, we consider 21 different OPE experiments in each setting.
-To provide precise performance assessments, in all the 21 experiments of both Logging 1 and 2 we generate $n_{data}=100$ realizations of the logging dataset, as done in the procedure outlined in [40].
-The performance metrics are subsequently averaged over all realizations to yield more robust performance estimates with their confidence intervals.
-Results
-The results obtained from Logging 1 are reported in Figure 8 while those of Logging 2 in Figure 8.
-Since these synthetic experiments are the same of [40], we use their optimal values for the other hyperparameters such as the network structure, regularization, learning rate.
-In both cases our proposed AutoOPE outperforms PAS-IF across almost all evaluation policies in terms of Spearman’s rank correlation coefficient.
-In terms of Relative Regret, AutoOPE performs similar to PAS-IF in the first half settings of the Logging 1 experiments.
-However, for high values of $\beta_{e}$, PAS-IF slightly outperform AutoOPE.
-In the Logging 2 set of experiments, we see that for all the $\beta_{e}\geq 0$ AutoOPE outperforms PAS-IF, but PAS-IF performs better in the settings with $\beta_{e}<0$.
-The fact that AutoOPE exhibits better Spearman’s Rank Correlation indicates that, while both methods predict similar values for the best estimator, AutoOPE is more consistent in ranking all of them.
-Appendix G Infrastructure and Computational Time Comparison
+Appendix I INFRASTRUCTURE AND COMPUTATIONAL TIME COMPARISON
 For the experiments conducted in this paper, we employed different instances from AWS EC2, to match different needs of resources (GPUs, CPUs) according to the different experiments.
 The Operating System Ubuntu 20.04 was selected for all the tests.
-We used the c6i.32xlarge instance with 128 cores and 256 GB RAM for the synthetic dataset generation explained in Appendix C.2.
+We used the c6i.32xlarge instance with 128 cores and 256 GB RAM for the synthetic dataset generation explained in Appendix D.2.
 This instance allows to terminate the whole generation process in about 35 hours.
@@
 The instance used in this case is the m6i.32xlarge instance, with 128 cores and 512 GB RAM.
-The whole process last approximately 20 hours.
+The whole process last approximately 5 days.
 Instead, the experiments described in Section 4 require some hardware accelerator to speed-up the training of the neural network used by PAS-IF.
@@
 The AutoOPE is tested on two conditions, with the model loaded from cache, or from disk.
-Appendix H Ground-Truth Evaluation Policy Value Estimation
-H.1 Open Bandit Dataset
+Appendix J GROUND-TRUTH EVALUATION AND POLICY VALUE ESTIMATION
+J.1 Open Bandit Dataset
 With Open Bandit Dataset, the estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$, can be computed using as ground-truth policy value estimate of $\pi_{e}$, the average of the counterfactual rewards of the evaluation dataset $\mathcal{D}_{gt}$ collected under the evaluation policy $\pi_{e}$.
@@
 The following expression describes the estimated ground-truth MSE of an estimator:
-$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{gt}}\cdot\sum_{r_{t}\in\mathcal{D}_{gt}}r_{t}\right)^{2}$ (5)
-H.2 UCI Datasets and CIFAR-10
-In the experiment reported in Appendix F.3 and in the experiments in Section 4.2.2, we were able to compute the policy value of $\pi_{e}$ in the same way, since we use classification datasets.
+$\text{MSE}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{gt}}\cdot\sum_{r_{t}\in\mathcal{D}_{gt}}r_{t}\right)^{2}$ (6)
+J.2 UCI Datasets and CIFAR-10
+In the experiment reported in Appendix G.1 and in the experiments in Section 4.2.2, we were able to compute the policy value of $\pi_{e}$ in the same way, since we use classification datasets.
 The policy value $\pi_{e}$ is computed as a true counterfactual expected reward on the logging dataset: in fact, the reward associated with each action and context in the logging data is deterministic (the reward is 1 for the correct class label/action relative to the context, and 0 for all other labels/actions).
 In this experiment the evaluation dataset can be defined as $\mathcal{D}_{gt}\coloneqq\left\{x_{t},\left\{r_{t,v}\right\}_{v}^{|\mathcal{A}|},\left\{\pi_{e}(a|x_{t})\right\}_{a\in\mathcal{A}}\right\}_{t}^{n_{b}}$.
-The estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with CIFAR-10 and UCI data (transformed in Contextual Bandit data as described in Appendix F.3 and in Section 4.2.2), is computed as described by the following expression:
-$MSE(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{b}}\cdot\sum_{t}^{n_{b}}\sum_{v}^{|\mathcal{A}|}\pi_{e}(a_{v}|x_{t})\cdot r_{t,v}\right)^{2}$ (6)
-H.3 Synthetic Data Experiments
+The estimated ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with CIFAR-10 and UCI data (transformed in Contextual Bandit data as described in Appendix G.1 and in Section 4.2.2), is computed as described by the following expression:
+$\text{MSE}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}))=\left(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}})-\frac{1}{n_{b}}\cdot\sum_{t}^{n_{b}}\sum_{v}^{|\mathcal{A}|}\pi_{e}(a_{v}|x_{t})\cdot r_{t,v}\right)^{2}$ (7)
+J.3 Synthetic Data Experiments
 The procedure to estimate the ground-truth MSE of an estimator $\hat{V}_{m}$ predicting the evaluation policy value of $\pi_{e}$ with synthetic data, leverages on the possibility to produce synthetically a really big number of rounds under $\pi_{e}$.
@@
 4:  end for
-5:  $MSE(\hat{V}_{m}(\pi_{e}))\leftarrow\frac{1}{n_{e}}\cdot\sum_{s}^{n_{e}}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(s)})-\frac{1}{n_{gt}}\cdot\sum_{t}^{n_{gt}}\sum_{a_{j}\in\mathcal{A}}f_{\pi_{e}}(x_{t},a_{j};s)\cdot q(x_{t},a_{j};s))^{2}$
-In Appendix F.4 are specified mostly of the input parameters for Logging 1 and 2 used in Algorithm 4, a part from $n_{e},\ n_{gt}$, set with values $n_{e}=100,\ n_{gt}=1,000,000$.
-Appendix I Broader impact statement
+5:  $\text{MSE}(\hat{V}_{m}(\pi_{e}))\leftarrow\frac{1}{n_{e}}\cdot\sum_{s}^{n_{e}}(\hat{V}_{m}(\pi_{e};\mathcal{T}_{\text{ope}}^{(s)})-\frac{1}{n_{gt}}\cdot\sum_{t}^{n_{gt}}\sum_{a_{j}\in\mathcal{A}}f_{\pi_{e}}(x_{t},a_{j};s)\cdot q(x_{t},a_{j};s))^{2}$
+In Appendix G.2 are specified mostly of the input parameters for Logging 1 and 2 used in Algorithm 4, a part from $n_{e},\ n_{gt}$, set with values $n_{e}=100,\ n_{gt}=1,000,000$.
 With an improved performance in Estimator Selection, we improve the quality of OPE.